### v3

# Global Convergence of Iteratively Reweighted Least Squares for Robust Subspace Recovery

Abstract
Robust subspace estimation is fundamental to many machine learning and data analysis tasks. Iteratively Reweighted Least Squares (IRLS) is an elegant and empirically effective approach to this problem, yet its theoretical properties remain poorly understood. This paper establishes that, under deterministic conditions, a variant of IRLS with dynamic smoothing regularization converges linearly to the underlying subspace from any initialization. We extend these guarantees to affine subspace estimation, a setting that lacks prior recovery theory. Additionally, we illustrate the practical benefits of IRLS through an application to low-dimensional neural network training. Our results provide the first global convergence guarantees for IRLS in robust subspace recovery and, more broadly, for nonconvex IRLS on a Riemannian manifold.
keywords:
Robust subspace recovery, iteratively reweighted least squares, dynamic smoothing, nonconvex optimization
journal: Applied and Computational Harmonic Analysis
[…]
1 Introduction
Many real-world datasets in computer vision, machine learning, and bioinformatics exhibit underlying low-dimensional structure that can be effectively modeled using subspaces (Basri and Jacobs, 2003; Hartley and Zisserman, 2003; Lerman et al., 2015; Ding et al., 2020; Li et al., 2022; Yu et al., 2024; Wang et al., 2024; Ben Arous et al., 2024). A variety of methods have been proposed to estimate such a subspace. The most prevalent of these is Principal Component Analysis (PCA) (Pearson, 1901; Hotelling, 1933; Jolliffe, 1986), which finds the directions of maximum variance within a dataset. However, PCA is well-documented to be sensitive to outliers.
These limitations spurred research into robust variants of PCA that can handle outliers; see Lerman and Maunu (2018b) for a comprehensive review. We focus specifically on robust subspace recovery (RSR), where the goal is to identify a low-dimensional subspace that explains a subset of the data points (inliers) while ignoring the effect of corrupted data (outliers). Unlike classical PCA, which assumes Gaussian-distributed noise, RSR explicitly models the data as a mixture of inliers that lie near or on a low-dimensional subspace and outliers that may be arbitrarily positioned.
Formally, we consider a dataset of $n$ observations in $\mathbb{R}^{D}$, denoted by the multiset $\mathcal{X}$. This multiset consists of an inlier multiset and an outlier multiset, $\mathcal{X}=\mathcal{X}_{\mathrm{in}}\cup\mathcal{X}_{\mathrm{out}}$, where the inliers lie on a $d$-dimensional subspace $L_{\star}$. For clarity, we focus on the noiseless inlier setting in this paper. The RSR problem seeks to recover $L_{\star}$ given only $\mathcal{X}$.
In the following, we focus on global convergence. That is, we say that an iterative algorithm globally recovers a point if it converges to that point from any initialization. To accomplish this, we revisit the IRLS method for RSR from Lerman and Maunu (2018a), known as the Fast Median Subspace (FMS) algorithm. FMS seeks to minimize a robust, nonconvex, least absolute deviations energy function over the Grassmannian manifold $\mathscr{G}(D,d)$, which is the set of $d$-dimensional linear subspaces in $\mathbb{R}^{D}$.
The focus of this paper is efficient solutions to the RSR problem when the subspace $L_{\star}$ and the dataset $\mathcal{X}$ satisfies certain conditions. We say that a subspace $L$ and dataset $\mathcal{X}$ satisfy deterministic conditions if there exist nonrandom functions $g_{i}$ such that $g_{i}(\mathcal{X}_{\mathrm{in}},\mathcal{X}_{\mathrm{out}},L)\geq 0$, $i=1,\dots,k$. These functions encode conditions that compare the statistics of the inliers and outliers, and we formulate our specific conditions later in Section 4.1. At a high level, such conditions end up being sufficient for solving the RSR problem efficiently, while in general it is NP-hard (Clarkson and Woodruff, 2015).
The rest of this section is organized as follows: Section 1.1 presents the primary contributions of our work. Then, Section 1.2 discusses the organization of the paper, and Section 1.3 introduces the necessary notation.
1.1 Contributions
The goals of our work are threefold: (i) to develop strong theoretical guarantees for FMS, (ii) to generalize these guarantees to affine subspaces, and (iii) to provide further practical motivation for FMS. We have made significant progress in all three directions. In particular, our key contributions are:
  1. 1.
    Under deterministic conditions on $\mathcal{X}$ alone, FMS with a dynamically decreasing smoothing linearly converges to $L_{\star}$ from arbitrary initialization. To the best of our knowledge, this is the first global convergence result for IRLS on a Riemannian manifold in a nonconvex setting. As a corollary, under the same deterministic conditions on $\mathcal{X}$ alone, the FMS algorithm with fixed regularization converges to a point close to $L_{\star}$ from arbitrary initialization.
  2. 2.
    We extend FMS to the estimation of affine subspaces. For this new algorithm, we establish local linear convergence under a modified set of deterministic conditions on $\mathcal{X}$ and the initialization.
  3. 3.
    We run numerical experiments that demonstrate the advantages of dynamic smoothing for robust subspace recovery in adversarial settings.
  4. 4.
    We demonstrate the practical utility of FMS through an application to low-dimensional neural network training.
These results address a long-standing question of providing broad theoretical guarantees that align with the strong empirical performance of FMS, while also enabling new applications in modern machine learning. They are likely to be of broad interest to the machine learning community. In particular, IRLS is a specific instance of a more general paradigm in optimization known as Majorize-Minimization (MM) algorithms. This class of algorithms is based on upper bounding a complex objective function with a surrogate function whose minimum is simpler to compute. One can generate a sequence whose objective values monotonically decrease by iteratively minimizing the surrogate. MM algorithms have been widely used in signal processing and machine learning (Wiesel and Zhang, 2015; Sun et al., 2016), and can provide insight into machine learning building blocks like transformers (Yang et al., 2022).
Furthermore, our algorithms are specific instances of Riemannian optimization algorithms, which operate when the constraint set is a Riemannian manifold. There has been broad interest in studying properties of these methods (Absil et al., 2007; Boumal, 2023) and their practical application, for examples, see Vandereycken (2013); Jaquier et al. (2020); Hu et al. (2020).
Finally, there has also been widespread interest in analyzing the convergence of nonconvex optimization algorithms, which is notoriously difficult (Danilova et al., 2022). Fortunately, in many settings, nonconvex problems exhibit benign properties (Li et al., 2019) that enable favorable algorithmic behavior (Maunu et al., 2019). Most past results rely on gradient-based optimization, and this work complements them by providing a global convergence result for a nonconvex IRLS algorithm on a Riemannian manifold under a deterministic condition.
1.2 Paper Organization
In Section 2, we rigorously define the robust subspace recovery problem, provide the necessary mathematical background, and review related work. In Section 3, we propose our dynamic smoothing scheme in the FMS algorithm and its variant for estimating affine subspaces. In Section 4, we establish the theoretical guarantees for the proposed method. In Section 5, we run simulations on synthetic data that demonstrate the competitiveness of FMS with dynamic smoothing, and we also present the effective application of FMS in dimension-reduced training of neural networks.
1.3 Notation
For a finite multiset $\mathcal{X}$, $|\mathcal{X}|$ denotes the number of points. We use bold uppercase letters for matrices and bold lowercase letters for column vectors. Let $\boldsymbol{I}_{k}$ denote the identity matrix in $\mathbb{R}^{k\times k}$, where if $k$ is obvious, we simply write $\boldsymbol{I}$. For $d\leq D$, $\mathcal{O}(D,d)$ is the set of semi-orthogonal matrices in $\mathbb{R}^{D\times d}$, i.e., $\boldsymbol{U}\in\mathcal{O}(D,d)$ if and only if $\boldsymbol{U}^{\top}\boldsymbol{U}=\boldsymbol{I}_{d}$. For a $d$-dimensional linear subspace $L$, we denote by $\boldsymbol{P}_{L}$ the $D\times D$ matrix representing the orthogonal projector onto $L$, and $\boldsymbol{Q}_{L}=\boldsymbol{P}_{L^{\perp}}$ the projection onto the orthogonal complement of $L$. Projection matrices are symmetric and satisfy $\boldsymbol{P}^{2}_{L}=\boldsymbol{P}_{L}$ and $\mathrm{im}(\boldsymbol{P}_{L})=L$, where $\mathrm{im}(\cdot)$ denotes the column space of a matrix. For a matrix $\boldsymbol{A}$, $\|\boldsymbol{A}\|_{2}$ denotes its operator or spectral norm, and $\|\boldsymbol{A}\|_{F}$ denotes its Frobenius norm. Further, $\sigma_{j}(\boldsymbol{A})$ will denote the $j$th singular value of $\boldsymbol{A}$. For our inlier-outlier dataset $\mathcal{X}$, we let $n=|\mathcal{X}|$, $n_{\mathrm{in}}=|\mathcal{X}_{\mathrm{in}}|$, and $n_{\mathrm{out}}=|\mathcal{X}_{\mathrm{out}}|$. Writing $f(x)\gtrsim g(x)$ means that $f(x)\geq cg(x)$ for some absolute constant $c$.
2 Problem Setup and Related Work
This section sets the stage for our algorithm and our theoretical results. In Section 2.1, we rigorously define the problem we seek to solve and provide the necessary mathematical background. After this, we review related work in Section 2.2.
2.1 Problem Formulation and Background
We briefly summarize the necessary background to understand the primary contributions of this work, building upon the setup provided in the introduction. For simplicity, we formulate the problem for linear subspaces. The treatment of affine subspaces is discussed in Section 3.2.
Let $\operatorname{dist}(\boldsymbol{x},L)=\|\boldsymbol{Q}_{L}\boldsymbol{x}\|$ be the orthogonal distance between $\boldsymbol{x}\in\mathcal{X}$ and $L\in\mathscr{G}(D,d)$, where $\boldsymbol{Q}_{L}$ is the orthogonal projection onto the orthogonal complement of $L$. This paper investigates an IRLS approach to a nonconvex least absolute deviations formulation for RSR (Ding et al., 2006; Zhang et al., 2009; Maunu et al., 2019). This formulation minimizes the sum of distances from the data points to any $d$-dimensional subspace in $\mathscr{G}(D,d)$:
$\hat{L}=\operatorname*{arg\;min}_{L\in\mathscr{G}(D,d)}F(L):=\sum_{\boldsymbol{x}\in\mathcal{X}}\operatorname{dist}(\boldsymbol{x},L),$ (1)
This is an optimization problem over the Grassmannian manifold (Edelman et al., 1998). In fact, if we minimize the sum of squared distances $\sum_{\boldsymbol{x}\in\mathcal{X}}\operatorname{dist}^{2}(\boldsymbol{x},L)$, then we obtain the principal subspace. By using the non-squared distance instead, our estimator is less prone to the impact of large outliers.
An IRLS method for (1) iteratively applies weighted PCA. More specifically, at iteration $k$ its unregularized version minimizes $\sum_{\boldsymbol{x}\in\mathcal{X}}w_{\boldsymbol{x}}\operatorname{dist}^{2}(\boldsymbol{x},L)$ with weights $w_{\boldsymbol{x}}=1/\operatorname{dist}(\boldsymbol{x},L^{(k)})$ to find the updated subspace $L^{(k+1)}$. Here, $L^{(k)}$ is the subspace estimated in the previous iteration. While this method is elegant and widely used, it lacks global convergence and exact recovery guarantees, and this paper aims to fill this gap. Nevertheless, our theoretical analysis requires some modification of this simple procedure (see discussion of dynamic smoothing in Section 3). This analysis advances the state of nonconvex IRLS for Riemannian manifolds, specifically the Grassmannian $\mathscr{G}(D,d)$. Furthermore, it is natural to extend (1) by replacing $\mathscr{G}(D,d)$ with the set of affine subspaces. This is the subject of our algorithmic development in Section 3.2.
The analysis in this paper depends on the principal angles and vectors between subspaces as follows: For any $L_{0},L_{1}\in\mathscr{G}(D,d)$, we denote their $d$ principal angles by $\{\theta_{j}(L_{0},L_{1})\}_{j=1}^{d}$ and let the respective principal vectors for $L_{0}$ and $L_{1}$ be $\{\boldsymbol{u}_{j}(L_{0},L_{1})\}_{j=1}^{d}$ and $\{\boldsymbol{v}_{j}(L_{0},L_{1})\}_{j=1}^{d}$. Given bases $\boldsymbol{W}_{0}\in\mathcal{O}(D,d)$ for $L_{0}$ and $\boldsymbol{W}_{1}\in\mathcal{O}(D,d)$ for $L_{1}$, one can compute the principal vectors and angles by the singular value decomposition $\boldsymbol{W}_{0}^{\top}\boldsymbol{W}_{1}=\boldsymbol{R}_{0}\cos(\boldsymbol{\Theta})\boldsymbol{R}_{1}^{\top}$, in which case $\boldsymbol{\Theta}$ contains the principal angles, $\boldsymbol{W}_{0}\boldsymbol{R}_{0}$ contains the vectors $\boldsymbol{u}_{j}$ and $\boldsymbol{W}_{1}\boldsymbol{R}_{1}$ contains the vectors $\boldsymbol{v}_{j}$. See Lerman and Zhang (2014, Section 3.2.1) for a more detailed treatment.
2.2 Review of Related Work
In the following, we separately discuss related work in robust subspace recovery, iterative reweighted least squares, and the least absolute deviations-based RSR problem (1).
2.2.1 Robust Subspace Recovery (RSR)
RSR has been extensively studied in various contexts, with methods falling into several key categories. For a comprehensive review, see Lerman and Maunu (2018b). The distinction between our problem and robust PCA (RPCA) (Candès et al., 2011) should be noted. While both deal with corrupted low-rank structures, RPCA assumes elementwise corruptions throughout the data matrix, whereas RSR assumes some samples are wholly corrupted. Consequently, algorithms designed for one problem typically do not perform well on the other.
A significant direction involves robust covariance estimation. Methods like Tyler’s M-estimator (TME) (Tyler, 1987; Zhang, 2016) aim to robustly estimate an underlying covariance matrix from which the principal subspace can be extracted. These approaches have strong theoretical foundations regarding breakdown points, though their guarantees for subspace recovery are more limited in natural special settings. Stronger local guarantees were obtained for a subspace-constrained TME variant (Yu et al., 2024; Lerman and Zhang, 2024).
Many approaches frame RSR as an outlier or inlier identification problem, where the goal is first to remove outliers or find a set of inliers before applying standard PCA. Examples include RANSAC-based methods (Fischler and Bolles, 1981; Arias-Castro and Wang, 2017; Maunu and Lerman, 2019) and more recent approaches like Coherence Pursuit (Rahmani and Atia, 2017) and Thresholding-based Outlier Pursuit Cherapanamjeri et al. (2017). While conceptually simple, these methods often require careful parameter tuning and may struggle with high outlier percentages and higher intrinsic subspace dimensions.
2.2.2 Least Absolute Deviations-based RSR
In this work, we focus on a specific approach to solving the RSR problem that relies on least absolute deviations, i.e., the objective function in (1). This has been a focus of a major line of work, where one replaces the squared Euclidean distance in PCA with the absolute distance to achieve robustness. While this problem is computationally hard (Hardt and Moitra, 2013; Clarkson and Woodruff, 2015), several methods have been proposed to approximately solve it. These include convex relaxations like Outlier Pursuit (Xu et al., 2012) and REAPER (Lerman et al., 2015), as well as nonconvex approaches like geodesic gradient descent (Maunu et al., 2019), which can be easily replaced with projected subgradient descent (PSGM) (Zhu et al., 2018). The problem may also be addressed using Riemannian ADMM (Li et al., 2024), a general framework for manifold optimization with nonsmooth objectives. In addition, PSGM and manifold proximal point algorithm (Chen et al., 2019) have also been proposed for Dual Principal Component Pursuit (DPCP), which can be considered as a special case of (1), where $d$ is replaced by DdD-d (Zhu et al., 2018). Indeed, the analysis of these methods is analogous to that of Maunu et al. (2019). We remark that the analysis pursued in Zhu et al. (2018) only holds for the special case d=D1d=D-1. It was later generalized for any $d$ in Ding et al. (2021), again with similar ideas to Maunu et al. (2019), whose result holds for any subspace dimension.
This work investigates the FMS method proposed by Lerman and Maunu (2018a), which solves (1) using IRLS. This method is popular due to its simplicity and strong empirical performance. More recently, FMS has also been applied to hierarchical representations (Ding et al., 2024).
However, the theoretical analysis of the FMS algorithm is somewhat limited, as it generally guarantees convergence only to stationary points. In contrast, Maunu et al. (2019) investigates the landscape of the objective function and obtains a theoretical understanding of its associated gradient descent algorithm. However, they only use this analysis to prove local convergence of gradient descent, which requires significant parameter tuning. Also, until now, it has been unclear how this analysis can extend to other optimization frameworks like IRLS, which we will discuss next.
2.2.3 Iteratively Reweighted Least Squares (IRLS)
The method of IRLS is used to solve specific optimization problems with objective functions that are sums of loss functions applied to residuals (Holland and Welsch, 1977). In each step the method solves a weighted least squares problem.
This idea has been applied in numerous problems, such as geometric median (Weiszfeld, 1937; Kuhn, 1973), robust regression (Holland and Welsch, 1977; Peng et al., 2024), M-estimator in robust statistics (Maronna et al., 2006), $\ell_{q}$ norm-based sparse recovery (Daubechies et al., 2010; Kümmerle et al., 2020), low-rank matrix recovery (Fornasier et al., 2011), synchronization (Chatterjee and Govindu, 2017), and more. In particular, IRLS algorithms have also been used to solve the convex relaxation of (1) in Zhang and Lerman (2014); Lerman et al. (2015).
Despite their strong empirical performance, straightforward IRLS algorithms generally lack strong theoretical guarantees. To illustrate this, for the geometric median problem (Weiszfeld, 1937; Beck and Sabach, 2015), which minimizes $F(\boldsymbol{x})=\sum_{i=1}^{n}\|\boldsymbol{x}-\boldsymbol{x}_{i}\|$, an IRLS method iteratively reweights the data points with weights $w_{i}^{(k)}={1}/{\|\boldsymbol{x}^{(k)}-\boldsymbol{x}_{i}\|}$. If $\boldsymbol{x}^{(k)}=\boldsymbol{x}_{1}$, then $\boldsymbol{x}^{(k+1)}=\boldsymbol{x}_{1}$, as the weight for $\boldsymbol{x}_{1}$ becomes infinite. Daubechies et al. (2010) indicate the same problem of infinite weight in IRLS for sparse recovery. To address this issue, a common strategy is to introduce a smoothing parameter that bounds the weights from above, like choosing $w_{i}^{(k)}=1/(\epsilon^{2}+\|\boldsymbol{x}_{i}-\boldsymbol{x}\|^{2})^{1/2}$ or $w_{i}=1/\max(\epsilon,\|\boldsymbol{x}_{i}-\boldsymbol{x}\|)$. This strategy is used heuristically in Zhang and Lerman (2014); Lerman et al. (2015); Lerman and Maunu (2018a). Still, as pointed out in Peng et al. (2024), this method only converges to an $\epsilon$-approximate solution, and no theory exists on its global convergence rate.
Researchers have used different insights to reach a consensus on dynamically updating the weights (Daubechies et al., 2010; Fornasier et al., 2011). Recent advances in IRLS (Kümmerle et al., 2020; Peng et al., 2024) suggest applying the “dynamic smoothing Parameter” strategy, which adaptively chooses the regularization parameter based on the current weights. Under this strategy, they prove that the IRLS algorithm converges linearly to the solution of the original problem. However, we note that such an analysis of IRLS has not been applied to manifold optimization problems such as RSR, and the only theoretical guarantee in nonconvex optimization provided in Peng et al. (2024) yields a local result in which the initialization must be very close to the solution.
Another work that examines dynamic smoothing on manifolds is Beck and Rosset (2023), where the authors prove a sublinear convergence rate for a gradient method to a stationary point.
3 FMS: Review and Extension to Affine Subspaces
In this section, we outline our algorithmic contributions to the RSR problem. In Section 3.1, we review the FMS algorithm and discuss the addition of dynamic smoothing. Then, in Section 3.2, we discuss a novel affine variant of the FMS algorithm.
3.1 The FMS Algorithm
Here, we study the iteratively reweighted least-squares method (IRLS) to solve (1), which alternates between finding the top eigenvectors of a weighted covariance and updating the weights. We recall that the unregularized IRLS method at the $k$-th iteration is the solution to the weighted squared problem of
$L^{(k+1)}=\operatorname*{arg\;min}_{L}\sum_{\boldsymbol{x}\in\mathcal{X}}w^{(k)}_{\boldsymbol{x}}\operatorname{dist}(\boldsymbol{x},L)^{2},\,\,\,\text{where $w^{(k)}_{\boldsymbol{x}}=\frac{1}{\operatorname{dist}(\boldsymbol{x},L^{(k)})}$}.$ (2)
This solution coincides with the standard PCA procedure:
$L^{(k+1)}=T(L^{(k)})=\text{span of top $d$ eigenvectors of $\sum_{\boldsymbol{x}\in\mathcal{X}}w_{\boldsymbol{x}}^{(k)}\boldsymbol{x}\boldsymbol{x}^{\top}$}.$ (3)
To put forward an IRLS algorithm that gives a re-weighting without infinite components in the weight, it is natural to regularize the weights to be $w^{(k)}_{\boldsymbol{x}}=1/\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})$. The FMS algorithm by Maunu et al. (2019) adopts a fixed regularization $\epsilon_{k}$ for all $k\geq 0$.
In this work, we propose to choose $\epsilon_{k}$ in each step adaptively by incorporating a “Dynamic Smoothing” procedure. Let $\gamma$ denote a hyperparameter, which can be related to the inlier percentage (see Section 4). We denote by $q_{\gamma}$ the $\gamma$-quantile of the set of distances of the data points from the subspace computed at that iteration. That is,
$q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L^{(k)})\}_{\boldsymbol{x}\in\mathcal{X}})=\max\left\{a\in\mathbb{R}:\ \frac{|\{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L^{(k)})\leq a\}|}{|\mathcal{X}|}\leq\gamma\right\}.$
We then set the dynamic smoothing parameter
$\epsilon_{k}=\min(\epsilon_{k-1},q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L^{(k)})\}_{\boldsymbol{x}\in\mathcal{X}})),{\text{ where }\epsilon_{-1}:=\infty}.$ (4)
This choice, inspired by Kümmerle et al. (2020) and generalized to the subspace recovery setting, ensures that $\epsilon_{k}$ forms a nonincreasing sequence while remaining sufficiently large relative to the distances ${\operatorname{dist}(\boldsymbol{x},L^{(k)})}_{\boldsymbol{x}\in\mathcal{X}}$. We summarize the regularized IRLS in Algorithm 1, which we refer to as Fast Median Subspace with Dynamic Smoothing (FMS-DS).
Dynamic smoothing allows for the parameter $\epsilon$ to get smaller along the iterations. This leads to a procedure that can recover a solution to the unregularized problem by solving a sequence of regularized problems. This is made rigorous in Section 4.
The FMS-DS algorithm requires two inputs: an initial guess at the subspace $L^{(0)}$, and a hyperparameter $\gamma$. This hyperparameter can be estimated using prior knowledge or cross-validation. Its choice for specific inlier-outlier models is discussed in Section 4.
Algorithm 1 FMS-DS
Input: Set of observations $\mathcal{X}\subset\mathbb{R}^{D}$, initial estimated subspace $L^{(0)}\subset\mathscr{G}(D,d)$, $\gamma$: hyperparameter.
Output: An estimate of the underlying subspace.
Steps:
1:
Let $k=0$.
2: Choose $\epsilon_{k}$ suitably from $L^{(k)}$ and $\mathcal{X}$ by (4).
3: $L^{(k+1)}=T_{\epsilon_{k}}(L^{(k)})$, where
$T_{\epsilon}(L)=\text{span of top $d$ eigenvectors of \ $\sum_{i=1}^{n}\frac{\boldsymbol{x}_{i}\boldsymbol{x}_{i}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x}_{i},L),\epsilon)}$}.$ (5)
4: $k=k+1$.
5: Repeat Steps 2-4 until convergence.
3.2 The Affine FMS Algorithm
Most existing works on RSR (see Section 2.2) focus on recovering a robust linear $d$-dimensional subspace in $\mathbb{R}^{D}$. While our main theoretical results concern this linear case, our objective function (1) and its IRLS algorithm naturally extend to the affine setting, allowing for robust local recovery of an affine $d$-dimensional subspace.
Here we use $\mathscr{A}(D,d)$ to denote the set of all $d$-dimensional affine subspaces in $\mathbb{R}^{D}$. We distinguish between a representative pair $A=(L,\boldsymbol{m})\in\mathscr{G}(D,d)\times\mathbb{R}^{D}$, and the affine subspace $A]=\{\boldsymbol{m}+\boldsymbol{x}:\boldsymbol{x}\in L\}$, which is the affine subspace formed by translating the $d$-dimensional linear subspace $L$ by $\boldsymbol{m}$. The representative pair is not unique, since $(L,\boldsymbol{m})$ and $(L,\boldsymbol{m}^{\prime})$ refer to the same affine subspace whenever $\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{P}_{L}\boldsymbol{\delta}$ for some $\boldsymbol{\delta}\in\mathbb{R}^{D}$. In fact, $[A]$ is an equivalence class of representative pairs. For any $\boldsymbol{x}\in\mathbb{R}^{D}$, the distance to $[A]$ is given by $\operatorname{dist}(\boldsymbol{x},[A])=\|\boldsymbol{Q}_{L}(\boldsymbol{x}-\boldsymbol{m})\|$, which is invariant under the choice of representative. Thus, we sometimes also write $\operatorname{dist}(\boldsymbol{x},A)=\|\boldsymbol{Q}_{L}(\boldsymbol{x}-\boldsymbol{m})\|$.
We consider the following natural generalization of (1):
$\hat{A}]=\operatorname*{arg\;min}_{[A]\in\mathscr{A}(D,d)}F^{(a)}([A]):=\sum_{\boldsymbol{x}\in\mathcal{X}}\operatorname{dist}(\boldsymbol{x},[A])=\sum_{\boldsymbol{x}\in\mathcal{X}}\|\boldsymbol{Q}_{L}(\boldsymbol{x}-\boldsymbol{m})\|.$ (6)
This problem is invariant to the choice of representation, and so we rewrite it in parametrized form as
$\hat{A}=\operatorname*{arg\;min}_{A\in\mathscr{G}(D,d)\times\mathbb{R}^{D}}F^{(a)}(A):=\sum_{\boldsymbol{x}\in\mathcal{X}}\operatorname{dist}(\boldsymbol{x},A)=\sum_{\boldsymbol{x}\in\mathcal{X}}\|\boldsymbol{Q}_{L}(\boldsymbol{x}-\boldsymbol{m})\|.$
To solve this problem, we generalize Algorithm 1 based on a straightforward generalization of the argument in Section 3.1. The affine IRLS method iterates by solving the weighted least squares problem at the $k$th iteration, where we replace distances to linear subspaces with distances to affine subspaces. The affine setting of the FMS algorithm can thus be written as
$A^{(k+1)}=T^{(a)}_{\epsilon_{k}}(A^{(k)}):=\operatorname*{arg\;min}_{A\in\mathscr{A}(D,d)}\sum_{\boldsymbol{x}\in\mathcal{X}}w^{(k)}_{\boldsymbol{x}}\operatorname{dist}(\boldsymbol{x},A)^{2},\,\,\,\text{where $w^{(k)}_{\boldsymbol{x}}=\frac{1}{\max(\operatorname{dist}(\boldsymbol{x},A^{(k)}),\epsilon_{k})}$}.$ (7)
The solution can be obtained from the standard PCA procedure applied to $\boldsymbol{x}\in\mathcal{X}$ with weights $1/{\operatorname{dist}(\boldsymbol{x},A^{(k)})}$. Therefore, $T^{(a)}_{\epsilon}(A)=(L,\boldsymbol{m})$ is such that $\boldsymbol{m}=\sum_{\boldsymbol{x}\in\mathcal{X}}w_{\boldsymbol{x}}\boldsymbol{x}/\sum_{\boldsymbol{x}\in\mathcal{X}}w_{\boldsymbol{x}}$, where $w_{\boldsymbol{x}}=1/\max(\operatorname{dist}(\boldsymbol{x},A),\epsilon)$, and $L$ is the span of top $d$ eigenvectors of $\sum_{\boldsymbol{x}\in\mathcal{X}}w_{\boldsymbol{x}}(\boldsymbol{x}-\boldsymbol{m})(\boldsymbol{x}-\boldsymbol{m})^{\top}$.
Similar to the linear case in (4), we choose the regularization parameters using dynamic smoothing. In this case, the update for $\epsilon$ becomes
$\epsilon_{k}=\min(\epsilon_{k-1},q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},A^{(k)})\}_{\boldsymbol{x}\in\mathcal{X}})).$ (8)
Again, $\gamma$ is a hyperparameter that we discuss more in our theoretical results. We refer to the resulting algorithm as Affine Fast Median Subspace with Dynamic Smoothing (AFMS-DS). The method with fixed regularization is just referred to as Affine FMS (AFMS).
Remark 1.
AFMS is rotation and translation equivariant, in the sense that if $T_{\epsilon}^{(a)}(A)=(L,\boldsymbol{m})$, then $T_{\epsilon}^{(a)}(\boldsymbol{R}A+\boldsymbol{a})=(\boldsymbol{R}L,\boldsymbol{R}\boldsymbol{m}+\boldsymbol{a})$ for any $\boldsymbol{R}^{\top}\boldsymbol{R}=\boldsymbol{I}_{D}$ and $\boldsymbol{a}\in\mathbb{R}^{D}$. Furthermore, FMS is a special case of AFMS with the mean parameter fixed to be zero.
4 Theoretical Analysis
This section establishes theoretical guarantees for the (A)FMS-DS algorithms by showing that the algorithms recover an underlying linear or affine subspace under specific inlier-outlier models. We emphasize that, in the case of FMS-DS, we achieve a global result, namely, the algorithm converges to the underlying subspace from any initialization. The main theorem for FMS-DS is presented in Section 4.1, with further discussion of this result presented in Sections 4.1.1 and 4.1.2. We then establish a local convergence theorem for AFMS-DS in Section 4.2.
4.1 Main Result for Linear Subspace Recovery
This section establishes a linear convergence guarantee for Algorithm 1. The main result, Theorem 1 shows that the algorithm linearly converges to an underlying subspace from any initialization if certain conditions on the data are satisfied.
To state our theorem, we first define a regularized objective $F_{\epsilon}(L)$:
$F_{\epsilon}(L)$ $=\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L)>\epsilon}\operatorname{dist}(\boldsymbol{x},L)+\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L)\leq\epsilon}\left(\frac{1}{2}\epsilon+\frac{\operatorname{dist}(\boldsymbol{x},L)^{2}}{2\epsilon}\right).$ (9)
This is a natural extension of (1) as we view it as the objective function for IRLS with a fixed regularization parameter (Zhang and Lerman, 2014; Lerman et al., 2015; Daubechies et al., 2010; Lerman and Maunu, 2018a; Kümmerle et al., 2020; Peng et al., 2024).
Next, we introduce the two assumptions needed for our main theorem; we discuss them further in Section 4.1.1. Both of these assumptions are restrictions on the inlier-outlier dataset $\mathcal{X}$. The first assumption requires that low-dimensional subspaces that are not $L_{\star}$ may not contain a significant number of points.
Assumption 1.
Let $\gamma_{\star}=|\mathcal{X}_{\mathrm{in}}|/|\mathcal{X}|$ be the percentage of inliers and $\gamma$ be the hyperparameter for FMS-DS. For all (d1d-1)-dimensional linear subspaces $L_{0}\subset L_{\star}$ and $d$-dimensional linear subspaces $L\neq L_{\star}$, $|\mathcal{X}\cap(L_{0}\cup L)|/|\mathcal{X}|<\gamma\leq\gamma_{\star}/2$.
Remark 2.
Assumption 1 can be weakened by instead assuming that $L=L_{\star}$ is the unique subspace such that $q_{\gamma}(\{\boldsymbol{x},L\}_{\boldsymbol{x}\in\mathcal{X}})=0$. However, in this case, we can only establish convergence to $L_{\star}$, and the linear convergence rate guarantee no longer holds.
The second assumption requires the definition of two summary statistics $S_{\mathrm{in}}$ and $S_{\mathrm{out}}$.
$S_{\mathrm{in}}$ $=\min_{\boldsymbol{u}\in L_{\star}:\|\boldsymbol{u}\|=1}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}|\boldsymbol{u}^{\top}\boldsymbol{x}|,\ S_{\mathrm{out}}=\max_{L\in\mathscr{G}(D,d)}\left\|\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\frac{\boldsymbol{P}_{L}\boldsymbol{x}\boldsymbol{x}^{\top}\boldsymbol{P}_{L^{\perp}}}{\operatorname{dist}(\boldsymbol{x},L)}\right\|_{2}.$ (10)
Heuristically, the statistic $S_{\mathrm{in}}$ measures how well-spread the inliers are on the underlying subspace, while the statistic $S_{\mathrm{out}}$ shows how well-aligned the outliers are along any $d$-dimensional subspace. Both statistics are somewhat natural. A large $S_{\mathrm{in}}$ prevents inliers from lying on or close to low-dimensional subspaces of $L_{\star}$. The statistic $S_{\mathrm{in}}$ is zero when there is a direction in $L_{\star}$ that is orthogonal to all inliers, and it is large when every direction in $L_{\star}$ has many inliers significantly correlated with it. On the other hand, having a small $S_{\mathrm{out}}$ is also natural, since it prevents outliers from being too low-dimensional. We remark that the $S_{\mathrm{in}}$ statistic is the same as the permeance statistic in Lerman et al. (2015), and the $S_{\mathrm{out}}$ statistic is a modified version of the alignment statistic of Maunu et al. (2019). In particular, Maunu et al. (2019) contains an additional projection onto $L$ within the norm, which would make the numerator of (10) $\boldsymbol{P}_{L}\boldsymbol{x}\boldsymbol{x}^{\top}\boldsymbol{P}_{L^{\perp}}$.
In line with the discussion in the previous paragraph, the second assumption ensures that the outlier statistic $S_{\mathrm{out}}$ is sufficiently smaller than the inlier statistic $S_{\mathrm{in}}$ and an additional regularized inlier statistic. This holds in a wide variety of cases, see the discussions in Maunu et al. (2019); Maunu and Lerman (2019) as well as Section 4.1.1.
Assumption 2.
For the inlier-outlier dataset $\mathcal{X}$, there exists a $\theta_{0}\in(0,\pi/2)$, the following deterministic condition is satisfied:
$\cos\theta_{0}S_{\mathrm{in}}\geq{3}\sqrt{d}S_{\mathrm{out}}.$ (11)
We also impose the following assumption, which is necessary to ensure that the iterates remain within a bounded neighborhood around $L_{\star}$ for $k\geq 1$. This assumption depends on a fixed parameter $\alpha_{0}$ in $(0,\gamma/\gamma_{\star})$. For simplicity, in the concrete examples presented in Section 4.1.1 we set $\alpha_{0}=0.5\cdot\gamma/\gamma_{\star}.$
Assumption 3.
There exists $\alpha_{0}\in(0,\gamma/\gamma_{\star})$ such that the following conditions hold for all subspaces $L\in\mathscr{G}(D,d)$:
  • Dominance of a quantile of the inlier distances over their mean: There exists $\beta_{1}>0$ such that
    $\beta_{1}\cdot q_{\alpha_{0}}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}})\geq\mathrm{mean}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}).$ (12)
  • Possible dominance of a quantile of the outlier distances over the mean of the inlier norms: There exists $\beta_{2}\geq 1$ such that if $\gamma\leq 1-(1-\alpha_{0})\gamma_{\star}$, then
    $\beta_{2}\cdot q_{\frac{\gamma-\alpha_{0}\gamma_{\star}}{1-\gamma_{\star}}}\left(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\right)\geq q_{\alpha_{0}}\left(\left\{\|\boldsymbol{x}\|\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right).$ (13)
    If $\gamma>1-(1-\alpha_{0})\gamma_{\star}$, then (13) cannot hold, since the quantile percentage is larger than 1. In this case, we still define $\beta_{2}=1$ and use this parameter in the next condition.
  • Dominance of the inliers over the outliers in a spectral sense: For $\theta_{0}$ defined in Assumption 2 and $\beta:={10\beta_{1}}{\beta_{2}}$,
    $\sin\theta_{0}\cdot\min_{\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}\,:\,|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|}\sigma_{d}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\boldsymbol{x}\boldsymbol{x}^{\top}\right)\geq 2\beta\cdot\sigma_{1}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\boldsymbol{x}\boldsymbol{x}^{\top}\right).$ (14)
Here, (12) implies that the inliers are well-distributed over $L_{\star}$, capturing a property analogous to a lower bound on $S_{\mathrm{in}}$. Equation (13) ensures that no subspace in $\mathscr{G}(D,d)$ contains a certain fraction of the outliers, capturing a property analogous to Assumption 1. Lastly, (14) guarantees a spectral dominance of the inliers over the outliers.
Remark 3.
The appearance of $\cos(\theta_{0})$ and $\sin(\theta_{0})$ in Assumptions 2 and 3 introduces a tradeoff in this parameter. Namely, this means that $\theta_{0}$ should be chosen not too close to 0 or $\pi/2$, since in (11) or (14) the left hand sides become zero. For simplicity, we set $\theta_{0}=\pi/4$ in our later analysis of these conditions in Section 4.1.1.
With these assumptions, we are now ready to present our main theorem. This theorem guarantees the global convergence of FMS-DS with a linear rate.
Theorem 1.
[Global Linear Convergence of FMS-DS] Let $\mathcal{X}$ be a dataset satisfying Assumptions 1, 2, and 3 for the given constant $\gamma>0$. Then, the sequence $L^{(k)}$ generated by Algorithm 1 with dynamic smoothing hyperparameter $\gamma$ converges to $L_{\star}$ in the sense that $\lim_{k\to\infty}\|\boldsymbol{P}_{L^{(k)}}-\boldsymbol{P}_{L_{\star}}\|_{2}\to 0$. In addition, the sequence $F_{\epsilon_{k}}(L^{(k)})$ is nonincreasing and there exists a $0<c<1$ such that
$F(L^{(k)})-F(L_{\star})\leq c^{k}(F(L^{(0)})-F(L_{\star})).$
To prove Theorem 1, we rely on several lemmas. To improve readability, the proof of these lemmas is deferred to Section 6.1.
Proof of Theorem 1.
We begin the proof with two lemmas. The first lemma is a simple monotonicity property on the regularized objective function $F_{\epsilon}$ defined in (9).
Lemma 1 (Properties of smoothed objective function).
For any $\epsilon_{\mathrm{in}}<\epsilon_{2}$, $F_{\epsilon_{\mathrm{in}}}(L)\leq F_{\epsilon_{2}}(L)$. In addition, $F_{0}(L)=F(L)$.
The second lemma shows that the smoothed objective value is nonincreasing over iterations of Algorithm 1, where the amount of decrease is quantified.
Lemma 2 (Decrease over iterations).
For all $k\geq 1$, we have
$F_{\epsilon_{k}}(L^{(k)})-F_{\epsilon_{k}}(L^{(k+1)})\geq\frac{\|\boldsymbol{P}_{L^{(k)}}\boldsymbol{S}_{L^{(k)},\epsilon_{k}}\boldsymbol{Q}_{L^{(k)}}\|_{F}^{2}}{2\|\boldsymbol{S}_{L^{(k)},\epsilon_{k}}\|},$ (15)
where
$\boldsymbol{S}_{L,\epsilon}=\sum_{\boldsymbol{x}\in\mathcal{X}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\epsilon,\operatorname{dist}(\boldsymbol{x},L))}.$ (16)
Using the fact that $\epsilon_{k}$ is nonincreasing, Lemmas 1 and  2 imply that the sequence $F_{\epsilon_{k}}(L^{(k)})$ are nonincreasing:
$F_{\epsilon_{k}}(L^{(k)})-F_{\epsilon_{k+1}}(L^{(k+1)})\geq F_{\epsilon_{k}}(L^{(k)})-F_{\epsilon_{k}}(L^{(k+1)})\geq 0.$ (17)
As a result, by basic monotonic convergence theory, the sequence $F_{\epsilon_{k}}(L^{(k)})$ must converge. Consequently, $F_{\epsilon_{k}}(L^{(k)})-F_{\epsilon_{k+1}}(L^{(k+1)})$ converges to zero.
The next two lemmas imply that the sequence $\epsilon_{k}$ converges to zero. The third lemma helps bound the right-hand side (RHS) of (15) by a quantity proportional to $\epsilon_{k}$, but depending also on $L$, whereas the fourth lemma removes the dependence on $L$. We note that the first lower bound in the next lemma is on the Frobenius norm of $\boldsymbol{P}_{L}\boldsymbol{S}_{L,\epsilon}\boldsymbol{Q}_{L}$, which is a Riemannian gradient of $F_{\epsilon}(L)$, and we thus view it as establishing nonzero gradient.
Lemma 3 (Nonzero gradient, local guarantee).
If $L\in G(D,d)$, $\epsilon>0$ and at least half of the inliers are of distance larger than $\epsilon$ from $L$, that is, $|\{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}:\operatorname{dist}(\boldsymbol{x},L)>\epsilon\}|\geq\frac{1}{2}|\mathcal{X}_{\mathrm{in}}|$, then
$\|\boldsymbol{P}_{L}\boldsymbol{S}_{L,\epsilon}\boldsymbol{Q}_{L}\|_{F}\geq\frac{\cos\theta_{1}(L,L_{\star})}{2\sqrt{d}}S_{\mathrm{in}}-S_{\mathrm{out}}\ \text{ and }\ \|\boldsymbol{S}_{L,\epsilon}\|\leq\frac{\|\boldsymbol{X}\|^{2}}{\epsilon}.$
The fourth lemma shows that under Assumption 3, all iterations of FMS lies in a neighborhood of the true subspace $L_{\star}$. The size of this neighborhood is controlled by $\theta_{0}$, which appears in Assumption 3.
Lemma 4 ($L^{(k)}$ is in a local neighborhood of $L_{\star}$).
Under Assumption 3, for all iterations $L^{(k)}$, $\theta_{1}(L^{(k)},L_{\star})\leq\theta_{0}$.
Combining (17), Lemmas 2-4, and the first inequality in (11), we have
$F_{\epsilon_{k}}(L^{(k)})-F_{\epsilon_{k+1}}(L^{(k+1)})\geq\frac{\|\boldsymbol{P}_{L^{(k)}}\boldsymbol{S}_{L^{(k)},\epsilon_{k}}\boldsymbol{Q}_{L^{(k)}}\|_{F}^{2}}{2\|\boldsymbol{S}_{L^{(k)},\epsilon_{k}}\|}\geq\epsilon_{k}\frac{\left(\frac{S_{\mathrm{in}}\cos\theta_{0}}{2\sqrt{d}}-S_{\mathrm{out}}\right)^{2}}{2\|\boldsymbol{X}\|^{2}}\geq\epsilon_{k}\frac{S_{\mathrm{out}}^{2}}{8\|\boldsymbol{X}\|^{2}}.$ (18)
We concluded from (17) that $F_{\epsilon_{k}}(L^{(k)})-F_{\epsilon_{k+1}}(L^{(k+1)})$ converges to zero. Thus (18) implies that $\epsilon_{k}$ converges to zero as well.
The definition (4) implies that there exists a subsequence $1\leq k_{1}\leq k_{2}\leq\cdots$ such that $\epsilon_{k_{l}}=q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L^{(k_{l})}\}_{\boldsymbol{x}\in\mathcal{X}})$. Since $\epsilon_{k}\to 0$, by Assumption 1, $\operatorname{dist}(\boldsymbol{x},L^{(k_{l})})\to 0$ for all inliers $\boldsymbol{x}$. This implies in particular that $\boldsymbol{u}_{j}(L^{(k)},L_{\star})\to L_{\star}$ for all $j=1,\dots,d$. Since $\operatorname{dist}(\boldsymbol{x},L^{(k)})^{2}=\sum_{j=1}^{d}\sin^{2}\theta_{j}(L^{(k)},L_{\star})|\boldsymbol{x}^{\top}\boldsymbol{u}_{j}(L^{(k)},L_{\star})|^{2}$, and by nonzero $S_{\mathrm{in}}$, we can choose a subset of inliers $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{d}$ that span $L_{\star}$ and we find that the sum of these distances, $\sum_{i=1}^{d}\sum_{j=1}^{d}\sin^{2}\theta_{j}(L^{(k)},L_{\star})|\boldsymbol{x}_{i}^{\top}\boldsymbol{u}_{j}(L^{(k)},L_{\star})|^{2}$, also converges to zero. For each $j$ there exists at least one $i$ such that $\boldsymbol{x}_{i}^{\top}\boldsymbol{u}_{j}(L^{(k)},L_{\star})$ is nonzero and bounded below. Thus, all principal angles must converge to zero, and as a result $\boldsymbol{P}_{L^{(k)}}-\boldsymbol{P}_{L_{\star}}$ converges to zero.
We now proceed with the proof of linear convergence. To establish that $F_{\epsilon_{k}}(L^{(k)})$ converges linearly, it suffices to show the existence of a constant $C_{0}>0$ such that
$F_{\epsilon_{k}}(L^{(k)})-F(L_{\star})<\epsilon_{k}C_{0}.$ (19)
In particular, we set
$C_{0}=\Bigg{(}\Big{(}1+\frac{\pi\sqrt{d}S_{\mathrm{out}}}{{2}S_{\mathrm{in}}}\Big{)}|\mathcal{X}_{\mathrm{in}}|/c_{3}+|\mathcal{X}|/2\Bigg{)},$ (20)
where $c_{3}$ will be specified in Lemma 6. The combination of (18) and (19) implies
$F_{\epsilon_{k}}(L^{(k)})-F_{\epsilon_{k+1}}(L^{(k+1)})\geq\frac{F_{\epsilon_{k}}(L^{(k)})-F(L_{\star})}{C_{0}}\cdot\frac{S_{\mathrm{out}}^{2}}{8\|\boldsymbol{X}\|^{2}}.$
As a result, $F_{\epsilon_{k}}(L^{(k)})$ converges linearly to $F(L_{\star})$.
We establish (19) through our final two lemmas. The next lemma establishes a bound on the ratio of outlier and inlier objective values based on $S_{\mathrm{in}}$ and $S_{\mathrm{out}}$. Let the inlier and outlier objectives be written as
$F_{\mathrm{in}}(L)=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\operatorname{dist}(\boldsymbol{x},L)\text{ and }F_{\mathrm{out}}(L)=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\operatorname{dist}(\boldsymbol{x},L).$
Lemma 5 (Bound on the objective value).
If the principal angles between $L$ and $L_{\star}$ are $\theta_{1},\cdots,\theta_{d}$ and $F_{\mathrm{in}}$ and $F_{\mathrm{out}}$ are as above, then
$\frac{|F_{\mathrm{out}}(L)-F_{\mathrm{out}}(L_{\star})|}{F_{\mathrm{in}}(L)-F_{\mathrm{in}}(L_{\star})}\leq\frac{\sqrt{d}S_{\mathrm{out}}\sum_{j=1}^{d}\theta_{j}}{S_{\mathrm{in}}\sum_{j=1}^{d}\sin\theta_{j}}.$ (21)
The final lemma establishes a connection between the quantiles of the set of distances and the inlier objective value that is useful for controlling the iterates $\epsilon_{k}$.
Lemma 6.
Under Assumption 1, there exists $c_{3}>0$ such that for all $L\in\mathscr{G}(D,d)$,
$q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}})\geq c_{3}\frac{F_{\mathrm{in}}(L)}{|\mathcal{X}_{\mathrm{in}}|}.$ (22)
To prove (19), we begin with the natural bound
$F_{\epsilon_{k}}(L^{(k)})$ $=\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L^{(k)})>\epsilon_{k}}\operatorname{dist}(\boldsymbol{x},L^{(k)})+\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L^{(k)})\leq\epsilon_{k}}\left(\frac{1}{2}\epsilon_{k}+\frac{\operatorname{dist}(\boldsymbol{x},L^{(k)})^{2}}{2\epsilon_{k}}\right).$
$\leq\sum_{\boldsymbol{x}\in\mathcal{X}}\operatorname{dist}(\boldsymbol{x},L^{(k)})+\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L^{(k)})\leq\epsilon_{k}}\epsilon_{k}$
$\leq F(L^{(k)})+\epsilon_{k}|\mathcal{X}|/2.$
Second, Lemma 5 implies
$F(L^{(k)})-F(L_{\star})$ $=F_{\mathrm{in}}(L^{(k)})-F_{\mathrm{in}}(L_{\star})+F_{\mathrm{out}}(L^{(k)})-F_{\mathrm{out}}(L_{\star})$
$\leq\left(1+\frac{S_{\mathrm{out}}\sum_{j=1}^{d}\theta_{j}}{\frac{\sum_{j=1}^{d}\sin\theta_{j}}{\sqrt{d}}S_{\mathrm{in}}}\right)\left(F_{\mathrm{in}}(L^{(k)})-F_{\mathrm{in}}(L_{\star})\right).$
Third, when $\epsilon_{k}=q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L^{(k)})\}_{\boldsymbol{x}\in\mathcal{X}})$, Lemma 6 implies
$F_{\mathrm{in}}(L^{(k)})-F_{\mathrm{in}}(L_{\star})=F_{\mathrm{in}}(L^{(k)})\leq|\mathcal{X}_{\mathrm{in}}|\epsilon_{k}/c_{3}.$
Combining the three equations above,
$F_{\epsilon_{k}}(L^{(k)})-F(L_{\star})$ $\leq(F(L^{(k)})-F(L_{\star}))+\epsilon_{k}|\mathcal{X}|/2$ (23)
$\leq\left(1+\frac{S_{\mathrm{out}}\sum_{j=1}^{d}\theta_{j}}{\frac{\sum_{j=1}^{d}\sin\theta_{j}}{\sqrt{d}}S_{\mathrm{in}}}\right)\left(F_{\mathrm{in}}(L^{(k)})-F_{\mathrm{in}}(L_{\star})\right)+\epsilon_{k}|\mathcal{X}|/2$
$\leq\left(\left(1+\frac{S_{\mathrm{out}}\sum_{j=1}^{d}\theta_{j}}{\frac{\sum_{j=1}^{d}\sin\theta_{j}}{\sqrt{d}}S_{\mathrm{in}}}\right)|\mathcal{X}_{\mathrm{in}}|/c_{3}+|\mathcal{X}|/2\right)\epsilon_{k}.$
Using $\theta\leq\frac{\pi}{2}\sin\theta$ for $\theta\in[0,\pi/2]$, and the definition of $C_{0}$ in (20),
$\left(\left(1+\frac{S_{\mathrm{out}}\sum_{j=1}^{d}\theta_{j}}{\frac{\sum_{j=1}^{d}\sin\theta_{j}}{\sqrt{d}}S_{\mathrm{in}}}\right)|\mathcal{X}_{\mathrm{in}}|/c_{3}+|\mathcal{X}|/2\right)\leq C_{0}$
and (19) holds. When $\epsilon_{k}\neq q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L^{(k)})\}_{\boldsymbol{x}\in\mathcal{X}})$, then there exists $l<k$ such that $\epsilon_{k}=\epsilon_{l}=q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L^{(l)})\}_{\boldsymbol{x}\in\mathcal{X}})$ and we have
$F_{\epsilon_{k}}(L^{(k)})\leq F_{\epsilon_{l}}(L^{(l)})\leq C_{0}\epsilon_{l}=C_{0}\epsilon_{k},$
where the second inequality follows from (23).
We conclude with a few remarks.
Remark 4 (Choice of $c_{3}$ in Lemma 6).
By the proof of Lemma 6, it is sufficient to find $c_{3}$ such that no more than $\gamma/2\gamma_{\star}$ percentage of the inliers are in the set
$\{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}:|\boldsymbol{u}_{1}^{\top}\boldsymbol{x}|\leq c_{3}\cdot\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\|\boldsymbol{x}\|\},$
for any $\|\boldsymbol{u}_{1}\|=1$, where $\mathrm{mean}$ is the average value of the set, and no more than $\gamma/2(1-\gamma_{\star})$ percentage of the outliers are contained in the region
$\{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}:|\boldsymbol{x}^{\top}\boldsymbol{u}_{2}|\leq c_{3}\cdot\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\|\boldsymbol{x}\|\}$
for any $\|\boldsymbol{u}_{2}\|=1$.
Remark 5.
Compared to recent guarantees for IRLS (Daubechies et al., 2010; Kümmerle et al., 2020; Peng et al., 2024), our work represents a step forward in two key ways. We analyze an IRLS algorithm on a Riemannian manifold, $\mathscr{G}(D,d)$, and establish global convergence for a nonconvex problem, both of which are novel results in the literature. In contrast, most analyses rely on convex optimization, ensuring global convergence. The only work on IRLS with nonconvex $\ell_{p}$ minimization ($p<1$) by Peng et al. (2024) achieves only local convergence.
4.1.1 Discussion of Assumptions
Assumption 1 ensures that $L_{\star}$ contains more points than all other low-dimensional subspaces by some margin. In the Generalized Haystack Model discussed below, it is satisfied with probability 1.
Assumptions 2 and 3 represent generic conditions on the inlier-outlier dataset, $\mathcal{X}$, that can be satisfied in a wide variety of settings. We can illustrate distinct examples where the conditions of these assumptions hold using data models inspired by Maunu et al. (2019); Maunu and Lerman (2019). To simplify our setting, we will assume Gaussian inliers and consider two models of outliers: Gaussian outliers and arbitrary outliers on the sphere. To further simplify, we set $\gamma=\gamma_{*}/2$, and we verify Assumptions 2 and 3 with $\theta_{0}=\pi/4$ and $\alpha_{0}=\gamma/(2\gamma_{\star})=1/4$.
The case of Gaussian inliers and Gaussian outliers is the Generalized Haystack Model (Maunu et al., 2019). In this model, we assume that $\boldsymbol{x}\in\mathcal{X}$ are i.i.d. sampled from a distribution $\mu=\alpha_{\mathrm{in}}\mu_{\mathrm{in}}+\alpha_{\mathrm{out}}\mu_{\mathrm{out}}$, that is, it is sampled from $\mu_{\mathrm{in}}$ with probability $\alpha_{\mathrm{in}}$ and sampled from $\mu_{\mathrm{out}}$ with probability $\alpha_{\mathrm{out}}$, where $\alpha_{\mathrm{in}}+\alpha_{\mathrm{out}}=1$. In particular, $\mu_{\mathrm{out}}$ is the distribution of inliers and is assumed to be $N(0,\boldsymbol{\Sigma}_{\mathrm{in}}/d)$, and $\mu_{\mathrm{in}}$ is the distribution of inliers and is assumed to be $N(0,\boldsymbol{\Sigma}_{\mathrm{out}}/D)$. In addition, we assume that $C\sigma_{\mathrm{in}}^{2}\boldsymbol{P}_{L_{\star}}/d\succeq\boldsymbol{\Sigma}_{\mathrm{in}}\succeq\sigma_{\mathrm{in}}^{2}\boldsymbol{P}_{L_{\star}}/d$ and $c\sigma_{\mathrm{out}}^{2}\boldsymbol{I}/D\preceq\boldsymbol{\Sigma}_{\mathrm{out}}\preceq\sigma_{\mathrm{out}}^{2}\boldsymbol{I}/D$, i.e., their condition numbers are $O(1)$. Under this model, we verify that Assumptions 2 and 3 hold, as stated in the following proposition.
Proposition 1 (Generalized Haystack Model).
Suppose that inliers and outliers follow the Generalized Haystack Model, $n\gtrsim D^{3}\log D$, $\gamma=\gamma_{\star}/2$, $\alpha_{0}=1/4$, $\theta_{0}=\pi/4$, and there exists an absolute constant $C$ such that $\gamma_{\star}\geq\min\left(\frac{C}{D-d},\,0.01\right)$. Then, Assumptions 2 and 3 hold with high probability as $n,D\to\infty$, provided that
$\frac{n_{\mathrm{in}}}{n_{\mathrm{out}}}\gtrsim\frac{\sigma_{\mathrm{out}}}{\sigma_{\mathrm{in}}}\cdot\frac{d}{\sqrt{D(D-d)}}+\frac{\sigma_{\mathrm{out}}^{2}}{\sigma_{\mathrm{in}}^{2}}\cdot\frac{d}{D}.$ (24)
For a proof, see Section 6.3. This can be compared with the bound for small sample sizes derived for a different nonconvex method in Maunu et al. (2019), which states that
$\frac{n_{\mathrm{in}}}{n_{\mathrm{out}}}\gtrsim\frac{\sigma_{\mathrm{out}}}{\sigma_{\mathrm{in}}}\cdot\frac{d}{\sqrt{D(D-d)}}.$
The two bounds are of the same order when $\sigma_{\mathrm{out}}\approx\sigma_{\mathrm{in}}$ or $\sigma_{\mathrm{out}}\leq\sigma_{\mathrm{in}}$, and they only differ by a factor of $\sqrt{\frac{\sigma_{\mathrm{in}}}{\sigma_{\mathrm{out}}}}\cdot\sqrt{\frac{D-d}{D}}$ when $\sigma_{\mathrm{out}}\geq\sigma_{\mathrm{in}}$. One could also consider an asymptotic regime as in Lerman and Maunu (2018a); Maunu et al. (2019), where it is possible to take the ratio $n_{\mathrm{in}}/n_{\mathrm{out}}$ to zero for fixed $d$ and $D$ as $n\to\infty$ in the Haystack model, where $\boldsymbol{\Sigma}_{\mathrm{in}}=\boldsymbol{P}_{L_{\star}}/d$ and $\boldsymbol{\Sigma}_{\mathrm{out}}=\boldsymbol{I}/D$.
In the case of adversarial outliers inspired by Maunu and Lerman (2019), we allow outliers to follow an arbitrary distribution on the sphere, and we assume that inliers follow a Gaussian distribution $N(\boldsymbol{0},\sigma_{\mathrm{in}}^{2}\boldsymbol{P}_{L_{\star}}/d)$. We refer to this as an adversarial model because outliers can follow any distribution. We restrict outliers to the sphere because our objective is not invariant to scale. We also assume that $\gamma_{\star}\geq 3/4$, so that in Assumption 3 we only need to show that (12) and (14) hold. The following proposition verifies Assumptions 2 and 3.
Proposition 2 (Adversarial Outliers).
Suppose that $\mathcal{X}\subset S^{D-1}$, $\gamma=\gamma_{\star}/2$, $\alpha_{0}=1/4$, $\theta_{0}=\pi/4$, and $\gamma_{\star}\geq 3/4$. Then, Assumptions 2 and 3 hold with high probability provided that
$\frac{n_{\mathrm{in}}}{n_{\mathrm{out}}}\gtrsim\frac{d}{\sigma_{\mathrm{in}}^{2}}+\frac{\sqrt{d}}{\sigma_{\mathrm{in}}}.$ (25)
The proof of this proposition is given in Section 6.4. We see that both Assumptions 2 and 3 are satisfied once $n_{\mathrm{in}}/n_{\mathrm{out}}=O(d)$. This matches the result of Maunu and Lerman (2019).
4.1.2 Extensions of Theory
There exist a couple of straightforward extensions of our theoretical results that we do not include in this paper.
Stability to noise: The FMS-DS algorithm exhibits empirical stability to noise. Our analysis can be extended to show that FMS-DS converges to a neighborhood of $L_{\star}$ since any $L$ outside this neighborhood can not be a fixed point. Technical analysis is deferred to future work.
Fixed smoothing parameters: The FMS algorithm (Maunu et al., 2019) uses the fixed value $\epsilon_{k}=\epsilon$ for $k\geq 1$. However, they are only able to guarantee convergence for the Haystack model. A corollary of Theorem 1 allows us to give a more general guarantee of linear convergence of FMS under our generic conditions. Its proof follows from the proof of Lemma 3, and it also implies that all stationary points of the regularized objective function $F_{\epsilon}(\boldsymbol{x})$ lie within a small neighborhood of $L_{\star}$.
Corollary 1 (Global convergence of FMS with fixed $\epsilon$).
Under the Assumptions 1, 2, and 3 and if $\epsilon_{k}=\epsilon$ for all $k\geq 1$, then in $2\big{\|}\sum_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{x}\boldsymbol{x}^{\top}\big{\|}\frac{F_{\epsilon}(L^{(0)})-F(L_{\star})}{\epsilon{S_{\mathrm{in}}\cos\theta_{0}-S_{\mathrm{out}}}}$ iterations, FMS reaches a subspace $L^{(k)}$ satisfying $q_{0.5}(\{\operatorname{dist}(\boldsymbol{x},L^{(k)})\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}))\leq\epsilon$.
4.2 Affine Subspace Recovery
This section presents a local theoretical guarantee for the AFMS-DS algorithm in Section 3.2. While weaker than Theorem 1 due to stricter assumptions, it is, to our knowledge, the first theoretical guarantee for robust affine subspace estimation.
We now present our assumptions for the main theorem on affine subspace recovery. First, we assume without loss of generality (WLOG) that $\boldsymbol{m}_{\star}=0$ due to the translation equivariance of the IRLS method. Second, the following assumption generalizes Assumption 1 to the affine setting.
Assumption 4.
For all (d1d-1)-dimensional affine subspaces $A_{0}]\subset[A_{\star}]$ and $d$-dimensional affine subspaces $A]\neq[A_{\star}]$ in $\mathscr{A}(D,d)$, $|\mathcal{X}\cap([A_{0}]\cup[A])|/|\mathcal{X}|<\gamma\leq\gamma_{\star}/2$.
Finally, we also have a generalization of our generic condition in Assumption 2 to the affine setting. To do this, we define a notion of distance between a representative $A=(L,\boldsymbol{m})\in\mathscr{G}(D,d)\times\mathbb{R}^{D}$ and an equivalence class $A^{\prime}]=[(L^{\prime},\boldsymbol{m}^{\prime})]\in\mathscr{A}(D,d)$:
$\operatorname{dist}(A,[A^{\prime}])=\sqrt{\sum_{i=1}^{d}\theta_{i}^{2}(L,L^{\prime})+\operatorname{dist}^{2}(\boldsymbol{m},[A^{\prime}])},$ (26)
where $\operatorname{dist}(\boldsymbol{m},[A^{\prime}])=\operatorname{dist}(\boldsymbol{m}-\boldsymbol{m}^{\prime},L^{\prime})$. We discuss this more in Section 4.2.1. Due to the translation equivariance of the AFMS method, without loss of generality, we can assume that $A_{\star}]$ is a linear subspace, so that $A_{\star}]=L_{\star}$. This simplification allows us to write $\operatorname{dist}(A_{\star},[A])=\operatorname{dist}((L_{\star},0),[A])$.
Assumption 5.
There exists $c_{0}<\pi/2$ such that for any $A]\in\mathscr{A}(D,d)$ satisfying $\operatorname{dist}(A_{\star},[A])\leq c_{0}$,
$\frac{1}{\cos^{2}c_{0}}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\operatorname{dist}(\boldsymbol{x},[A])$ $\geq\max 4\Big{(}\frac{\pi}{2}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\operatorname{dist}(\boldsymbol{P}_{L_{\star}}\boldsymbol{x},[A]),\theta_{1}(L,L_{\star})^{2}\Big{\|}P_{L_{\star}^{\perp}}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\boldsymbol{x}\boldsymbol{x}^{\top}P_{L_{\star}^{\perp}}\Big{\|}\Big{)}.$ (27)
In Assumption 5, (27) implies that the influence of the inliers is larger than that of the outliers. Compared with the assumptions for Theorem 1, these assumptions are more restrictive.
Our main theoretical result on AFMS-DS guarantees that it converges locally and can be viewed as a generalization of Theorem 1.
Theorem 2.
[Local convergence of AFMS-DS] Under Assumptions 4 and 5 and $\operatorname{dist}(A_{\star},[A^{(k)}])\leq c_{0}$ for all $k\geq 1$, where $A^{(k)}$ is the sequence generated by AFMS-DS, then $A^{(k)}]$ converges to $A_{\star}]$. In addition, the sequence $F_{\epsilon_{k}}(A^{(k)})$ is nonincreasing and converges to $F(A_{\star})$ linearly.
The proof is a natural extension of the argument presented in Theorem 1 and is provided in Section 6.2. However, in the affine setting, Assumption 5 becomes more challenging to simplify to a more interpretable form. We conjecture that further analysis may yield results analogous to Assumptions 2 and 3, potentially leading to global convergence guarantees. We leave this investigation for future work.
Our result in Theorem 2 relies on an additional condition versus Theorem 1, namely that $\operatorname{dist}(A_{\star},[A^{(k)}])\leq c_{0}$ for all $k\geq 1$, but it does not have an analog of Assumption 3. We note, however, that by following the proof of Theorem 2, if the initialization is chosen such that $F_{\epsilon_{\mathrm{out}}}(A^{(0)})$ is sufficiently small (see definition of $F_{\epsilon}(A)$ in (53)), then this condition can be satisfied due to monotonicity of $F_{\epsilon_{k}}(A^{(k)})$. In particular, monotonicity of this sequence implies that the sequence $A^{(k)}$ is constrained to lie in a level set of $F_{\epsilon_{\mathrm{out}}}$. Consequently, a “good initialization” assumption is sufficient for the convergence of AFMS-DS.
4.2.1 Discussion of Assumption 4
Following the discussion in Section 4.1.1, we analyze Assumption 4 under both the adversarial and Haystack models. Unlike the earlier spherical setting, we do not consider spherized models here, as they are not well-suited for affine subspace recovery.
Instead, we adopt a probabilistic framework in which a fraction $\alpha_{\mathrm{in}}$ of the samples are inliers and $\alpha_{\mathrm{out}}=1-\alpha_{\mathrm{in}}$ are outliers. Inliers are drawn from the distribution $\mu_{\mathrm{in}}=\mathcal{N}(\boldsymbol{m}_{\star},\boldsymbol{\Sigma}_{\mathrm{in}}/d)$, where $\boldsymbol{\Sigma}_{\mathrm{in}}$ is a positive semidefinite matrix with range in $L_{\star}$ and $\sigma_{d}(\boldsymbol{\Sigma}_{\mathrm{in}})\geq 1$. Outliers are either drawn from a general distribution $\mu_{\mathrm{out}}$ (affine adversarial model), or from a Gaussian $\mathcal{N}(\boldsymbol{m}_{\star},\boldsymbol{\Sigma}_{\mathrm{out}}/D)$ with $\boldsymbol{\Sigma}_{\mathrm{out}}\preceq\sigma_{\mathrm{out}}^{2}\boldsymbol{I}$ (Affine Haystack model). As in the proof of our main theorem, we can assume without loss of generality that $\boldsymbol{m}_{\star}=\boldsymbol{0}$ and identify $A_{\star}]$ with $L_{\star}$.
Proposition 3 (Affine Adversarial Model).
Under the affine adversarial model, Assumption 4 holds if
$\frac{\alpha_{\mathrm{in}}}{\alpha_{\mathrm{out}}}>O\left(\max\left({\sqrt{d}}{\int_{\boldsymbol{x}\sim\mu_{\mathrm{out}}}\|P_{L_{\star}}\boldsymbol{x}\|},\ 1,\ c_{0}\cos^{2}(c_{0})\sqrt{d}\left\|\int_{\boldsymbol{x}\sim\mu_{\mathrm{out}}}P_{L_{\star}^{\perp}}\boldsymbol{x}\boldsymbol{x}^{\top}P_{L_{\star}^{\perp}}\right\|\right)\right).$
Proposition 3 suggests that if $c_{0}$ is chosen sufficiently small and the distribution $\mu$ has support of magnitude $O(1)$, then the outlier ratio must be bounded by $O(1/\sqrt{d})$. This is better than the linear case, where the outlier ratio is bounded by $O(1/d)$, because the analysis of the affine case does not use the extra Assumption 3; instead, we assume that the initialization is sufficiently good, specifically by requiring that $c_{0}$ is small enough.
Proposition 4 (Affine Haystack Model).
Under the adversarial Haystack model, and assuming $\sigma_{\mathrm{out}}\leq\sqrt{D/d}$, Assumption 4 and equation (27) hold provided that
$\frac{\alpha_{\mathrm{in}}}{\alpha_{\mathrm{out}}}\geq\max\left(1,\,O\left(c_{0}{\frac{\sqrt{d}}{D}}\sigma_{\mathrm{out}}^{2}\right)\right).$
We note that the required inlier fraction for the Affine Haystack model is $O(1)$, in contrast to the $O(d/\sqrt{D(D-d)})$ requirement in the Generalized Haystack Model discussed in Section 4.1.1. However, the Affine Haystack model allows outliers to have magnitudes up to a factor of $\sqrt{D/d}$ larger.
The $O(1)$ inlier requirement is intuitive. For instance, in the case of estimating a 0-dimensional affine subspace (i.e., computing the median), the breakdown point is $0.5$, meaning that at least half of the data must be inliers for robust recovery. However, since the outliers in the Affine Haystack model are relatively benign, we expect the breakdown point of AFMS to be lower. We leave a precise characterization of this threshold to future work.
Discussion of Distance: For our analysis, we assume that the set of inliers $\mathcal{X}_{\mathrm{in}}$ lies on a $d$-dimensional affine subspace $A_{\star}]=[(L_{\star},\boldsymbol{m}_{\star})]$. This defines a distance between a representative for $[A]$ and the affine subspace $A^{\prime}]$. To our knowledge, this notion of distance is new in the literature and may be of independent interest. This is not a true metric on the affine Grassmannian since it depends on the choice of $\boldsymbol{m}$ in the representation for $[A]$. Note, however, that $\operatorname{dist}(A,[A^{\prime}])=0$ if and only if $A]=[A^{\prime}]$.
5 Numerical Simulations
In this section, we present experiments on synthetic and real data. Our first experiment in Section 5.1 compares the FMS algorithm with various competitive RSR methods. The second and third experiments in Sections 5.2 and 5.3 investigate the performance of FMS under various regularization strategies, highlighting the effectiveness of the proposed dynamic smoothing approach in escaping saddle points. The final experiment in Section 5.4 demonstrates the practical utility of FMS dimensionality reduction in training neural networks. The code used for our experiments and implementations is available at: https://github.com/alp-del/Robust-PCA.
5.1 Experiment 1: Subspace Recovery Across Methods
For the first experiment, we evaluate the performance of four competitive RSR algorithms on synthetic datasets generated under a semi-adversarial inlier-outlier model. The algorithms chosen are RANSAC (Maunu and Lerman, 2019), STE (Yu et al., 2024), TME (Tyler, 1987; Zhang, 2016), and FMS. We also include PCA as a baseline subspace recovery method. Comparisons with other RSR algorithms have been conducted in prior work and consistently show inferior performance across a range of regimes (see, e.g., Lerman and Maunu (2018b)), especially when one is interested in exact recovery. Although RANSAC and TME were not originally developed for RSR, we use their RSR-adapted versions in our experiments. Finally, we note that the most common implementation of DPCP (Zhu et al., 2018) is equivalent to FMS, but targets the orthogonal complement of the desired subspace. As a result, it yields the same solution as FMS when using the same initialization and parameters.
The inlier-outlier models used here aim to demonstrate the performance of these RSR algorithms in a semi-adversarial setting. The datasets consist of two components: inliers lying on a subspace of dimension $d\in\{3,10\}$ and sampled from a standard Gaussian distribution supported on that subspace, and outliers lying on a separate subspace of dimension $d_{\mathrm{out}}\in\{1,5,10\}$ and independently sampled from a standard Gaussian distribution supported on the outlier subspace. WLOG, we assume that the ambient dimension is $D=d+d_{\mathrm{out}}$. As a final step, we normalize all points so that they lie on the unit sphere in $\mathbb{R}^{D}$ to remove effects of scale.
The goal is to compare the performance of the RSR algorithms in these semi-adversarial settings. Therefore, we fix the total number of data points at $n=160$, and we vary the proportion of inliers and outliers to see where each algorithm breaks down.
Since we are comparing iterative algorithms, we ensure a fair comparison of running times by setting the maximum number of iterations to 200 for the RANSAC, FMS, TME, and STE methods. In particular, we emphasize that RANSAC is executed with only 200 candidate samples. This choice is motivated by the fact that the computational complexity of a single iteration of FMS or STE is comparable to that of evaluating one candidate sample in RANSAC.
For the STE method, we use a gamma parameter of 2, and for TME we use regularization of $\epsilon=10^{-10}$. We run FMS with fixed $\epsilon=10^{-10}$, as well as dynamic smoothing with two choices of $\gamma=0.5$ and $\gamma=0.1$. To ensure statistical reliability, we generate 200 datasets for each pair $d$ and $d_{\mathrm{out}}$, and the average performance is reported. FMS is initialized using the PCA subspace, and TME and STE are initialized with the sample covariance.
Let the orthonormal matrix $\boldsymbol{P}\in\mathbb{R}^{D\times d}$ represent the basis of the true inlier subspace, and let $\widetilde{\boldsymbol{P}}\in\mathbb{R}^{D\times d}$ denote the basis of the estimated subspace. We measure the subspace estimation error by $\left\|\widetilde{\boldsymbol{P}}\widetilde{\boldsymbol{P}}^{\top}-\boldsymbol{P}\boldsymbol{P}^{\top}\right\|_{2},$ which corresponds to the sine of the largest principal angle between the true subspace and its estimate. Due to the fact that we are in a noiseless regime, we plot all errors on a log scale. Since error is reported as the geometric mean of the errors over the 200 datasets, the result is the average log-error.
$d=3$ $d=10$
$d_{\mathrm{out}}=1$
Refer to caption Refer to caption
$d_{\mathrm{out}}=5$
Refer to caption Refer to caption
$d_{\mathrm{out}}=10$
Refer to caption Refer to caption
Figure 1: Performance comparison of PCA, TME, RANSAC, STE, and FMS on synthetic data, reporting the geometric mean of the error over 200 repetitions on a log scale. This corresponds to the averaged log-error. We vary outlier dimension across rows and inlier dimension across columns as follows: top row: outlier dimension $d_{\mathrm{out}}=1$, middle row: $d_{\mathrm{out}}=5$, bottom row: $d_{\mathrm{out}}=10$. Left column: inlier dimension $d=3$, right column: $d=10$. As we can see, FMS-DS with small $\gamma$ performs well across a range of settings. RANSAC performs well for small $d$, but fails for larger $d$ due to the fact that its runtime is exponential in $d$, and we cap the number of iterations.
As illustrated in Figure 1, FMS has a comparable performance as STE and TME and usually outperforms them, particularly when the inlier dimension is higher. RANSAC, which follows a fundamentally different approach, performs well in low-dimensional settings but degrades significantly as dimensionality increases. This is because RANSAC estimates candidate subspaces from randomly selected sets of $d$ samples, and its success depends on all $d$ samples being inliers. This condition becomes exponentially unlikely as $d$ grows larger.
Since STE, FMS, and TME are applied with fixed regularization parameters, as no dynamic smoothing schemes are available for them, they do not reach machine precision and have performance comparable with FMS with fixed regularization $\epsilon=10^{-10}$. On the other hand, applying FMS with fixed $\epsilon=10^{-15}$, which we do not display to prevent clutter, has the same performance as FMS-DS with $\gamma=0.1$, and thus it can reach machine precision.
5.2 Experiment 2: Effect of Regularization in FMS
We investigate how different regularization strategies influence the performance of FMS in a setting where the initialization is chosen adversarially. In particular, we aim to demonstrate that dynamic smoothing can avoid saddle points that FMS with fixed regularization may become stuck at.
We compare the following variants:
  • FMS with fixed regularization parameters of $\epsilon=10^{-3},10^{-10}$, and $10^{-15}$.
  • FMS-DS where $\epsilon$ is chosen by (4) with $\gamma=0.1$ and $0.5$. Here, we set the initial value $\epsilon^{(0)}$ to be the $\gamma$-quantile of the $\{\operatorname{dist}(\boldsymbol{x},L_{0})\}_{\boldsymbol{x}\in\mathcal{X}}$, where $L_{0}$ represents the initial subspace.
Similar to Experiment 1, the synthetic dataset is constructed with an inlier subspace dimension of $d=3$ and an outlier dimension of $d_{\mathrm{out}}=1$. The ambient space has dimension $D=d+d_{\mathrm{out}}=4$. The dataset contains a total of 200 data points. To demonstrate that the dynamic smoothing strategy can escape local minima or stationary points, we initialize the algorithm at a stationary point of the FMS objective—specifically, a subspace with two directions in the inlier subspace and one direction orthogonal to it. We again generate 200 datasets in this model and average the results for each algorithm. In this case, FMS is initialized with two directions within the inlier subspace and one direction orthogonal to it that contains the outliers.
Figures 2 and 3 report the results of this experiment. In Figure 2, we report the averaged log-error of the different methods as a function of outlier percentage. We also report the failure rate, which corresponds to FMS getting stuck at a bad stationary point, as a function of outlier percentage in Figure 3. As shown in these figures, FMS-DS with larger $\gamma$ significantly outperforms both fixed regularization approaches, regardless of the fixed regularization strength.
It is worth noting that FMS with a large regularization parameter consistently results in higher error due to the overly strong regularization. On the other hand, while FMS with small regularization can perform well in some cases, it begins to fail when the inlier percentage is $\geq 12\%$, in part because it becomes trapped in stationary points. In contrast, FMS with dynamic smoothing achieves consistently better performance across a range of outlier ratios.
Figure 2: Averaged log-error versus outlier percentage for the experiment with adversarial initialization. Here, the inlier subspace dimension is 3 and the outlier subspace is a 1-dimensional subspace orthogonal to it. FMS is initialized with two directions within the inlier subspace and one direction orthogonal to it.
Full View
Zoomed-in View
Figure 3: Failure rate of the various FMS algorithms under the same setting as Figure 2, for different outlier percentages. Left: full range of outlier percentages. Right: zoomed-in view.
5.3 Experiment 3: Iteration-Wise Error Analysis
We further investigate the convergence behavior of FMS by plotting the estimation error over iterations for each regularization strategy. We follow the settings of $d$ and $d_{\mathrm{out}}$ from Experiment 2, but fix the number of inliers at 100 and outliers at 30. The algorithm is initialized using the PCA subspace as in Experiment 1, or the orthogonal subspace as in Experiment 2.
In Figure 4, we plot the log-error versus iteration for FMS with the different regularization schemes. The left display of Figure 4 uses PCA initialization and shows that it is comparable to FMS with a fixed small regularization of $\epsilon=10^{-15}$. The rate of convergence is slightly slower than FMS with small regularization ($\epsilon=10^{-15}$). In contrast, for FMS with larger fixed regularization, the estimation error plateaus once it reaches the same order of magnitude as the regularization parameter. The right display of Figure 4 demonstrates the performance of the various regularization strategies with the poor initialization of Experiment 2. As we can see, FMS-DS with $\gamma=0.5$ and FMS with fixed $\epsilon=10^{-3}$ are more effective at escaping the saddle point. However, the fixed regularization is not able to adapt and plateaus at the fixed error of $\approx 10^{-3}$ while FMS-DS converges to the global solution. FMS-DS with $\gamma=0.1$ and FMS with small fixed $\epsilon=10^{-15}$ perform comparably.
PCA Initialization (Exp. 1)
Adversarial Initialization (Exp. 2)
Figure 4: Averaged log-error versus iteration for FMS with different regularization strategies. On the left, we demonstrate convergence with PCA initialization as in Experiment 1, and on the right, we demonstrate convergence with the orthogonal initialization of Experiment 2. The left plot demonstrates that in settings with good initialization and lack of bad stationary points, FMS with small fixed $\epsilon=10^{-15}$ and FMS-DS with small $\gamma=0.1$ both perform well. On the other hand, for bad initialization, we see that FMS with larger $\epsilon$ and FMS-DS with larger $\gamma$ escape the saddle point faster.
5.4 A Practical Demonstration of the FMS Algorithm
While the previous section provides a new theoretical analysis of the FMS algorithm and extends it to affine settings, this analysis applies only in inlier-outlier settings. However, it is typically unknown a priori whether a dataset satisfies this condition. To further motivate the FMS algorithm and its role in robust subspace estimation, we present an application related to low-dimensional neural network training.
A line of work studies the use of a subspace-constrained optimization. One example is neural network training, where Li et al. (2022) use Dynamic Linear Dimensionality Reduction (DLDR) to show that training in a PCA subspace can improve generalization. Unfortunately, this procedure of subspace estimation lies at odds with the current understanding of neural network training. We argue that using vanilla PCA may be misguided since recent evidence indicates that stochastic gradient descent in neural network training exhibits heavy-tailed noise (Simsekli et al., 2019; Zhou et al., 2020).
The implementation details of DLDR are as follows (Li et al., 2022). First, during training with SGD, we sample $t$ steps of neural network parameters $\{\boldsymbol{w}_{1},\boldsymbol{w}_{2},\dots,\boldsymbol{w}_{t}\}\subset\mathbb{R}^{D}$. The weights are then centered by subtracting the mean, and the PCA subspace is computed. The network is then retrained, with gradients projected to the PCA subspace.
Motivated by this, we replace the PCA subspace in the previous procedure with FMS and AFMS. For completeness, we also report the application of TME, PCA, and spherical PCA (SPCA), which computes the PCA subspace of the normalized dataset $\boldsymbol{x}/|\boldsymbol{x}|:\boldsymbol{x}\in\mathcal{X}$} (Locantore et al., 1999). We do not report dynamic smoothing, as it did not change the performance of FMS. In particular, dynamic smoothing typically affects the algorithm’s rate of convergence and precision, but neither is essential in this example.
Specifically, we trained CIFAR-10 (Krizhevsky, 2009) on ResNet-20, CIFAR-100 (Krizhevsky, 2009) on ResNet-32, and Tiny-ImageNet (Deng et al., 2009) on ResNet-18. The CIFAR-10 dataset consists of 60,000 color images of size 32×32 pixels, categorized into 10 classes. It contains 50,000 training images and 10,000 test images. The CIFAR-100 dataset is an extension of CIFAR-10, containing 100 classes instead of 10. Each class has 600 images, making a total of 60,000 images. The dataset consists of 50,000 training images and 10,000 test images of size 32×32 pixels. The Tiny ImageNet dataset contains 200 classes with images resized to 64×64 pixels. It includes 100,000 training images and 10,000 validation images.
For all three data sets, the images were normalized using their channel-wise means and variances. The deep neural networks were trained with the SGD optimizer using a weight decay of 1e-4, a momentum of 0.9, and a batch size of 128. The learning rate was initially set to 0.1 and decayed to 0.01 over 100 epochs. Model parameters were sampled at the end of every epoch, after which FMS, PCA, and TME were applied to estimate the subspace. When training within this subspace using projected-SGD, we maintained the same batch size and momentum as in the original SGD setup. The learning rate was set to 1 initially and decayed to 0.1 over 30 epochs.
To assess robustness, we introduced additional label corruption by randomly selecting a fraction of the training data and assigning random labels to them. The testing accuracies of our simulations are visualized in Figure 5. The first row illustrates the accuracies of all methods applied to three datasets reduced to 20 dimensions, while the second row shows the accuracies when 15% of the labels are randomly corrupted. Note that AFMS and FMS have the same performance in all scenarios. PCA and TME also perform similarly. As shown in Figure 5, label corruption significantly degrades the performance of SGD. In contrast, training in low-dimensional subspaces demonstrates that FMS outperforms the other methods.
CIFAR-10
Refer to captionRefer to caption
CIFAR-100
Refer to captionRefer to caption
Tiny Imagenet
Refer to captionRefer to caption
Figure 5: Comparison of PCA, SPCA, TME, FMS, and AFMS for low-dimensional neural network training using three different datasets (left: CIFAR-10, center: CIFAR-100, right: Tiny ImageNet). Top row: subspace dimension fixed at 20 and varying corruption, bottom row: corruption level fixed at 15% and varying subspace dimension. For sufficiently high dimensions, the subspace recovered by (A)FMS outperforms those obtained by other RSR methods.
6 Additional Proofs
In this section, we provide additional proofs that were deferred for readability. Section 6.1 contains the proofs of Lemmas 16, which were stated during the proof of Theorem 1. Section 6.2 presents the proof of Theorem 2, which builds upon the ideas used in the proof of Theorem 1. Lastly, Section 6.3 provides the proof of Proposition 1, and Section 6.5 provides the proofs of Propositions 3 and 4.
6.1 Proof of Supplementary Lemmas for Theorem 1
Proof of Lemma 1.
Assume that $\epsilon_{\mathrm{in}}<\epsilon_{2}$. We remind ourselves that
$F_{\epsilon}(L)$ $=\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L)>\epsilon}\operatorname{dist}(\boldsymbol{x},L)+\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L)\leq\epsilon}\left(\frac{1}{2}\epsilon+\frac{\operatorname{dist}(\boldsymbol{x},L)^{2}}{2\epsilon}\right).$
We proceed by cases for each $\boldsymbol{x}$.
If $\operatorname{dist}(\boldsymbol{x},L)>\epsilon_{2}$, then the corresponding term in $F_{\epsilon_{\mathrm{in}}}$ is equal to the term in $F_{\epsilon_{2}}$. Similarly, if $\operatorname{dist}(\boldsymbol{x},L)\leq\epsilon_{\mathrm{in}}$, then the term is again equal. It remains to study the case where $\epsilon_{\mathrm{in}}<\operatorname{dist}(\boldsymbol{x},L)\leq\epsilon_{2}$. The result follows from the fact that
$|z|\leq\frac{\epsilon}{2}+\frac{z^{2}}{2\epsilon}$
for all $\epsilon>0$. In particular, WLOG assume $z>0$, and the inequality is equivalent to
$z^{2}+\epsilon^{2}-2z\epsilon=(z-\epsilon)^{2}\geq 0.$
Proof of Lemma 2.
Before proving Lemma 2, we prove the following lemma on the quadratic growth of the trace function $f:\mathscr{G}(D,d)\rightarrow\mathbb{R}$ defined by $f(L)=\mathrm{tr}(\boldsymbol{P}_{L}\boldsymbol{\Sigma})$.
Lemma 7.
For a positive semi-definite matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{D\times D}$ and $\hat{L}=\operatorname*{arg\;max}_{L}f(L)$,
$f(\hat{L})-f(L)\geq\frac{\|\boldsymbol{P}_{L}\boldsymbol{\Sigma}\boldsymbol{P}_{L^{\perp}}\|_{F}^{2}}{\|\boldsymbol{\Sigma}\|}.$
Proof of Lemma 7.
In the proof we denote
$\boldsymbol{\Sigma}]_{L_{1},L_{2}}=\boldsymbol{U}_{L_{1}}^{\top}\boldsymbol{\Sigma}\boldsymbol{U}_{L_{2}}.$
and thus write
$\boldsymbol{\Sigma}=\begin{pmatrix}[\boldsymbol{\Sigma}]_{L,L}&[\boldsymbol{\Sigma}]_{L,L^{\perp}}\\ [\boldsymbol{\Sigma}]_{L^{\perp},L}&[\boldsymbol{\Sigma}]_{L^{\perp},L^{\perp}}\end{pmatrix}$. (28)
Inspired by Lerman and Zhang (2024), we define
$\boldsymbol{\Sigma}_{*}:=\begin{pmatrix}[\boldsymbol{\Sigma}]_{L,L}&[\boldsymbol{\Sigma}]_{L,L^{\perp}}\\ [\boldsymbol{\Sigma}]_{L^{\perp},L}&[\boldsymbol{\Sigma}]_{L^{\perp},L}([\boldsymbol{\Sigma}]_{L,L})^{-1}[\boldsymbol{\Sigma}]_{L,L^{\perp}}\end{pmatrix}$.
We note that $\boldsymbol{\Sigma}-\boldsymbol{\Sigma}_{*}$ is positive semi-definite. Indeed, three of its blocks are zero and the bottom-right block is $\boldsymbol{\Sigma}]_{L^{\perp},L^{\perp}}-[\boldsymbol{\Sigma}]_{L^{\perp},L}([\boldsymbol{\Sigma}]_{L,L})^{-1}[\boldsymbol{\Sigma}]_{L,L^{\perp}}$, which is positive semi-definite (Bhatia, 2009, Theorem 1.3.3). Since f(L^)f($\hat${L}) is the sum of the top $d$ eigenvalues of $\boldsymbol{\Sigma}$,
$f(\hat{L})=\sum_{i=1}^{d}\lambda_{i}(\boldsymbol{\Sigma})\geq\sum_{i=1}^{d}\lambda_{i}(\boldsymbol{\Sigma}_{*})=\mathrm{tr}(\boldsymbol{\Sigma}_{*})=\mathrm{tr}([\boldsymbol{\Sigma}]_{L,L})+\mathrm{tr}([\boldsymbol{\Sigma}]_{L^{\perp},L}([\boldsymbol{\Sigma}]_{L,L})^{-1}[\boldsymbol{\Sigma}]_{L,L^{\perp}}),$ (29)
where the first inequality follows from the fact that $\boldsymbol{\Sigma}-\boldsymbol{\Sigma}_{*}$ is positive semidefinite and the second equality follows from the fact that $\operatorname{rank}(\boldsymbol{\Sigma}_{*})\leq d$, as
$\boldsymbol{\Sigma}_{*}=\begin{pmatrix}[\boldsymbol{\Sigma}]_{L,L}\\ [\boldsymbol{\Sigma}]_{L^{\perp},L}\end{pmatrix}\begin{pmatrix}\boldsymbol{I}&([\boldsymbol{\Sigma}]_{L,L})^{-1}[\boldsymbol{\Sigma}]_{L,L^{\perp}}.\end{pmatrix}$
In addition,
$\mathrm{tr}([\boldsymbol{\Sigma}]_{L^{\perp},L}([\boldsymbol{\Sigma}]_{L,L})^{-1}[\boldsymbol{\Sigma}]_{L,L^{\perp}})$ $=\|[\boldsymbol{\Sigma}]_{L^{\perp},L}([\boldsymbol{\Sigma}]_{L,L})^{-1/2}\|_{F}^{2}$
$\geq\|[\boldsymbol{\Sigma}]_{L^{\perp},L}\|_{F}^{2}/\|[\boldsymbol{\Sigma}]_{L,L}\|$
$\geq\|[\boldsymbol{\Sigma}]_{L^{\perp},L}\|_{F}^{2}/\|\boldsymbol{\Sigma}\|.$
Recalling that $f(L)=\mathrm{tr}([\boldsymbol{\Sigma}]_{L,L})$, the lemma is proved. ∎
We now proceed with the proof of Lemma 2. We use an auxiliary function defined as follows:
$G_{\epsilon}(L,L_{0})=\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L_{0})>\epsilon}\Big{(}\frac{1}{2}\operatorname{dist}(\boldsymbol{x}_{i},L_{0})+\frac{\operatorname{dist}(\boldsymbol{x}_{i},L)^{2}}{2\operatorname{dist}(\boldsymbol{x}_{i},L_{0})}\Big{)}\\ +\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L_{0})\leq\epsilon}\Big{(}\frac{1}{2}\epsilon+\frac{\operatorname{dist}(\boldsymbol{x}_{i},L)^{2}}{2\epsilon}\Big{)}.$ (30)
Our proof is then based on Lemma 1, Lemma 7, and the following observations (see Section 5.1 of Lerman and Maunu (2018a) for proof):
$F_{\epsilon}(L)=G_{\epsilon}(L,L)$ (31)
$F_{\epsilon}(L_{0})\leq G_{\epsilon}(L,L_{0})$ with equality holds only when $L=L_{0}$
$L^{(k+1)}=\operatorname*{arg\;min}_{L}G_{\epsilon_{k}}(L,L^{(k)})$
From the observations in (31) we have
$F_{\epsilon_{k+1}}(L^{(k+1)})\leq F_{\epsilon_{k}}(L^{(k+1)})\leq G_{\epsilon_{k}}(L^{(k+1)},L^{(k)})\leq G_{\epsilon_{k}}(L^{(k)},L^{(k)})=F_{\epsilon_{k}}(L^{(k)}).$
This implies that
$F_{\epsilon_{k}}(L^{(k)})-F_{\epsilon_{k+1}}(L^{(k+1)})\geq G_{\epsilon_{k}}(L^{(k)},L^{(k)})-G_{\epsilon_{k}}(L^{(k+1)},L^{(k)}).$ (32)
Now, using the fact that
$\operatorname{dist}^{2}(\boldsymbol{x}_{i},L)^{2}=\mathrm{tr}(\boldsymbol{P}_{L^{\perp}}\boldsymbol{x}_{i}\boldsymbol{x}_{i}^{\top}\boldsymbol{P}_{L^{\perp}}),$
we can rewrite
$G_{\epsilon_{k}}(L,L_{0})$ $=\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L_{0})>\epsilon_{k}}\left(\frac{1}{2}\operatorname{dist}(\boldsymbol{x}_{i},L_{0})+\frac{\operatorname{dist}(\boldsymbol{x}_{i},L)^{2}}{2\operatorname{dist}(\boldsymbol{x}_{i},L_{0})}\right)\ +$
$\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L_{0})\leq\epsilon_{k}}\left(\frac{1}{2}\epsilon_{k}+\frac{\operatorname{dist}(\boldsymbol{x}_{i},L)^{2}}{2\epsilon}\right)$
$=\mathrm{tr}\left(\sum_{i}\frac{\boldsymbol{P}_{L^{\perp}}\boldsymbol{x}_{i}\boldsymbol{x}_{i}^{\top}\boldsymbol{P}_{L^{\perp}}}{2\max(\operatorname{dist}(x_{i},L_{0}),\epsilon)}\right)+C(L_{0})$
$=\frac{1}{2}\mathrm{tr}(\boldsymbol{P}_{L^{\perp}}\boldsymbol{S}_{L_{0},\epsilon_{k}}\boldsymbol{P}_{L^{\perp}})+C(L_{0})$
$=\frac{1}{2}\Big{[}\mathrm{tr}(\boldsymbol{S}_{L_{0},\epsilon_{k}})-\mathrm{tr}(\boldsymbol{P}_{L}\boldsymbol{S}_{L_{0},\epsilon_{k}}\boldsymbol{P}_{L})\Big{]}+C(L_{0})$
for some term that only depends on $L_{0}$. Notice that the third line implies that
$G_{\epsilon_{k}}(L,L_{0})=\frac{1}{2}\mathrm{tr}(\boldsymbol{S}_{L_{0},\epsilon_{k}}\boldsymbol{P}_{L^{\perp}})+C(L_{0}),$
and so we can apply Lemma 7 to find
$G(L^{(k)},L^{(k)})-G_{\epsilon_{k}}(L^{(k+1)},L^{(k)})\geq\frac{1}{2}\frac{\|\boldsymbol{P}_{L^{(k)}}\boldsymbol{S}_{L^{(k)},\epsilon_{k}}\boldsymbol{P}_{L^{(k)^{\perp}}}\|_{F}^{2}}{\|\boldsymbol{S}_{L^{(k)},\epsilon_{k}}\|}.$
Lemma 2 follows by the previous display and (31).∎
Proof of Lemma 3.
In this proof, we let $\boldsymbol{S}_{L,\epsilon}=\boldsymbol{S}_{L,\epsilon,in}+\boldsymbol{S}_{L,\epsilon,out}$, where
$\boldsymbol{S}_{L,\epsilon,in}=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L),\epsilon)},\,\,\boldsymbol{S}_{L,\epsilon,out}=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L),\epsilon)},$
and we will investigate $\boldsymbol{P}_{L}\boldsymbol{S}_{L,\epsilon,in}\boldsymbol{P}_{L^{\perp}}$ and $\boldsymbol{P}_{L}\boldsymbol{S}_{L,\epsilon,out}\boldsymbol{P}_{L^{\perp}}$ separately.
Let the principal vectors between $L$ and $L_{\star}$ be $\{\boldsymbol{u}_{j}\}_{j=1}^{d}$ and $\{\boldsymbol{v}_{j}\}_{j=1}^{d}$, corresponding to $L$ and $L_{\star}$, respectively, and let $\{\boldsymbol{w}_{j}\}_{j=1}^{d}$ be unit vectors in $L^{\perp}$ such that $\boldsymbol{v}_{j}=\cos\theta_{j}\,\boldsymbol{u}_{j}+\sin\theta_{j}\,\boldsymbol{w}_{j}$. Then we will show that there exists $1\leq j\leq d$ such that
$\boldsymbol{u}_{j}^{\top}\boldsymbol{S}_{L,\epsilon,in}\boldsymbol{w}_{j}>\frac{\cos\theta_{1}}{2\sqrt{d}}S_{\mathrm{in}}\ \text{ and }\ |\boldsymbol{u}_{j}^{\top}\boldsymbol{S}_{L,\epsilon,out}\boldsymbol{w}_{j}|<S_{\mathrm{out}}.$
We begin by analyzing $\boldsymbol{S}_{L,\epsilon,out}$. By the definition of $S_{\mathrm{out}}$,
$\|\boldsymbol{P}_{L}\boldsymbol{S}_{L,\epsilon,out}\boldsymbol{P}_{L^{\perp}}\|<S_{\mathrm{out}}$
and as a result, for all $1\leq j\leq d$,
$|\boldsymbol{u}_{j}^{\top}\boldsymbol{S}_{L,\epsilon,out}\boldsymbol{w}_{j}|<S_{\mathrm{out}}.$ (33)
We now analyze $\boldsymbol{P}_{L}\boldsymbol{S}_{L,\epsilon,in}\boldsymbol{P}_{L^{\perp}}$. We first investigate $\boldsymbol{u}_{j}^{\top}\boldsymbol{x}\boldsymbol{x}^{\top}\boldsymbol{w}_{j}$ for any $\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}$,
$\boldsymbol{u}_{j}^{\top}\boldsymbol{x}=\boldsymbol{u}_{j}^{\top}\left(\sum_{i=1}^{d}\boldsymbol{v}_{i}\boldsymbol{v}_{i}^{\top}\right)\boldsymbol{x}=\cos\theta_{j}\boldsymbol{v}_{j}^{\top}\boldsymbol{x},\ \boldsymbol{w}_{j}^{\top}\boldsymbol{x}=\boldsymbol{w}_{j}^{\top}\left(\sum_{i=1}^{d}\boldsymbol{v}_{i}\boldsymbol{v}_{i}^{\top}\right)\boldsymbol{x}=\sin\theta_{j}\boldsymbol{v}_{j}^{\top}\boldsymbol{x}.$ (34)
Next, we note that $P_{L^{\perp}}\boldsymbol{x}=\sum_{j=1}^{d}\boldsymbol{w}_{j}\boldsymbol{w}_{j}^{\top}\boldsymbol{x}$, and thus
$\operatorname{dist}(\boldsymbol{x},L)=\|P_{L^{\perp}}\boldsymbol{x}\|=\sqrt{\sum_{j=1}^{d}(\boldsymbol{w}_{j}^{\top}\boldsymbol{x})^{2}}=\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}(\boldsymbol{v}_{j}^{\top}\boldsymbol{x})^{2}}.$ (35)
In order to lower bound $\operatorname{dist}(\boldsymbol{x},L)$, we note that $\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}c_{j}}$ is a concave function of the variables $\sin^{2}\theta_{j}$. Consequently, using the result of Section 7.2.3 of Lerman et al. (2015)
$\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\operatorname{dist}(\boldsymbol{x},L)$ $=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}\cdot(\boldsymbol{v}_{j}^{\top}\boldsymbol{x})^{2}}\geq S_{\mathrm{in}}\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}}.$ (36)
Next, let us analyze $\sum_{j=1}^{d}\tan(\theta_{j})\,\boldsymbol{u}_{j}^{\top}\boldsymbol{S}_{L,\epsilon,in}\boldsymbol{w}_{j}.$ Applying (34) and (35), for an inlier $\boldsymbol{x}$, we have that
$\sum_{j=1}^{d}\tan\theta_{j}\frac{\sin\theta_{j}\cos\theta_{j}(\boldsymbol{v}_{j}^{\top}\boldsymbol{x})^{2}}{\max(\operatorname{dist}(\boldsymbol{x},L),\epsilon)}=\frac{\sum_{j=1}^{d}\sin^{2}\theta_{j}\cdot(\boldsymbol{v}_{j}^{\top}\boldsymbol{x})^{2}}{\max(\operatorname{dist}(\boldsymbol{x},L),\epsilon)}=\frac{\operatorname{dist}^{2}(\boldsymbol{x},L)}{\max(\operatorname{dist}(\boldsymbol{x},L),\epsilon)}.$
Recall that $\epsilon$ is chosen such that for at least half of the inliers, $\operatorname{dist}(\boldsymbol{x},L)\geq\epsilon$. Using this and (36),
$\sum_{j=1}^{d}\tan(\theta_{j})\boldsymbol{u}_{j}\boldsymbol{S}_{L,\epsilon,in}\boldsymbol{w}_{j}$ $=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\frac{\operatorname{dist}^{2}(\boldsymbol{x},L)}{\max(\operatorname{dist}(\boldsymbol{x},L),\epsilon)}$ (37)
$\geq\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}:\operatorname{dist}(\boldsymbol{x},L)>\epsilon}\operatorname{dist}(\boldsymbol{x},L)$
$\geq\frac{1}{2}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\operatorname{dist}(\boldsymbol{x},L)$
$\geq\frac{1}{2}S_{\mathrm{in}}\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}}.$
On the other hand,
$\frac{\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}}}{\sum_{j=1}^{d}\tan\theta_{j}}\geq\frac{\frac{1}{\sqrt{d}}{\sum_{j=1}^{d}\sin\theta_{j}}}{\frac{1}{\cos\theta_{1}}\sum_{j=1}^{d}\sin\theta_{j}}=\frac{\cos\theta_{1}}{\sqrt{d}}.$ (38)
As a result, (37) implies that
$\max_{j=1,\cdots,d}\boldsymbol{u}_{j}^{\top}\boldsymbol{S}_{L,\epsilon,in}\boldsymbol{w}_{j}\geq\frac{S_{\mathrm{in}}\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}}}{\sum_{j=1}^{d}\tan\theta_{j}}\geq\frac{\cos\theta_{1}}{2\sqrt{d}}S_{\mathrm{in}}.$
Denoting the index that achieves the maximum above by $j_{0}$ and using (33) and the above equation, the proof of Lemma 3 is concluded:
$\|\boldsymbol{P}_{L}\boldsymbol{S}_{L,\epsilon}\boldsymbol{P}_{L^{\perp}}\|_{F}\geq\boldsymbol{u}_{j_{0}}^{\top}\boldsymbol{S}_{L,\epsilon}\boldsymbol{w}_{j_{0}}\geq\boldsymbol{u}_{j_{0}}^{\top}\boldsymbol{S}_{L,\epsilon,in}\boldsymbol{w}_{j_{0}}+\boldsymbol{u}_{j_{0}}^{\top}\boldsymbol{S}_{L,\epsilon,out}\boldsymbol{w}_{j_{0}}\geq\frac{\cos\theta_{1}}{2\sqrt{d}}S_{\mathrm{in}}-S_{\mathrm{out}}.$
Proof of Lemma 4.
We will prove Lemma 4 by showing that for all $k\geq 0$, $\theta_{1}(L^{(k+1)},L_{\star})\leq\theta_{0}$.
We begin by the naive bound (recall that $\alpha_{0}\leq\gamma/\gamma_{\star}\leq 1/2$)
$q_{\alpha_{0}}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}})\leq\mathrm{median}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}})\leq 2\mathrm{mean}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}).$ (39)
The second bound follows from the fact that for positive numbers, the median is bounded by twice the mean. For any subspaces $L,L^{\prime}\in\mathscr{G}(D,d)$ such that $F_{\mathrm{in}}(L^{\prime})\geq F_{\mathrm{in}}(L)/5$, that is, $\mathrm{mean}(\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}})\geq\mathrm{mean}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}})/5$, (12) and (39) imply
$10\beta_{1}q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right)\geq q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L)\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right).$
The assumption (13) implies that
$q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right)$ $\leq q_{\alpha_{0}}\left(\left\{\|\boldsymbol{x}\|\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right)$
$\leq\beta_{2}\cdot q_{\frac{\gamma-\alpha_{0}\gamma_{\star}}{1-\gamma_{\star}}}\left(\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\right).$
Using this we find that
$q_{\gamma}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}}\right)$ $\geq\min\left(q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right),q_{\frac{\gamma-\alpha_{0}\gamma_{\star}}{(1-\gamma_{\star})}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\right)\right)$
$\geq\min\left(q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right),\frac{1}{\beta_{2}}q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right)\right)$
$\geq\frac{1}{\beta_{2}}q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right).$
We remark that when $\gamma\geq 1-(1-\alpha_{0})\gamma_{\star}$, then $(\gamma-\alpha_{0}\gamma_{\star})/(1-\gamma_{\star})\geq 1$, and hence, by setting $\beta_{2}=1$, the above condition is trivially satisfied under our assumption. In summary, we obtain the bound
$q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L)\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right)\leq{10}{\beta_{1}}\cdot q_{\alpha_{0}}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right)\leq\beta\cdot q_{\gamma}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\right\}_{\boldsymbol{x}\in\mathcal{X}}\right).$ (40)
Next, we show that at least $\alpha_{0}$ percentage of the inliers satisfy $\operatorname{dist}(\boldsymbol{x},L^{(k)})\leq\beta\epsilon_{k}$, i.e.,
$\mathcal{X}_{\mathrm{in}}^{\prime}=\left\{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}:\operatorname{dist}(\boldsymbol{x},L^{(k)})\leq\beta\epsilon_{k}\right\}\implies|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|.$ (41)
By the definition of $\epsilon_{k}$ in (4), there exists $l\leq k$ such that $\epsilon_{k}=q_{\gamma}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{(l)})\right\}_{\boldsymbol{x}\in\mathcal{X}}\right).$ We now consider two cases:
Case 1: $k=l$. In this case, (41) follows directly from (40) by taking $L=L^{\prime}=L^{(k)}$.
Case 2: $k>l$. Then we have $\epsilon_{l}>\epsilon_{l+1}=\cdots=\epsilon_{k}$. Since the objective function is non-increasing, we have
$F_{\epsilon_{k}}(L^{(l)})\geq F_{\epsilon_{k}}(L^{(k)})\geq F(L^{(k)}).$
Moreover, by the choice of $\epsilon_{k}$ and using $\gamma\leq\gamma_{\star}/2$, we get $F_{\epsilon_{k}}(L^{(l)})-F(L^{(l)})\leq\epsilon_{k}\gamma|\mathcal{X}|\leq F_{\mathrm{in}}(L^{(l)}),$ where the second inequality uses the fact that
$\epsilon_{k}=q_{\gamma}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{(l)})\right\}_{\boldsymbol{x}\in\mathcal{X}}\right)\leq q_{0.5}\left(\left\{\operatorname{dist}(\boldsymbol{x},L^{(l)})\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right),$
and again, for positive numbers, the median is bounded by twice the mean.
Hence,
$F_{\mathrm{in}}(L^{(l)})+F(L^{(l)})\geq F(L^{(k)}),\,\,\text{i.e.,}\,\,2F_{\mathrm{in}}(L^{(l)})+F_{\mathrm{out}}(L^{(l)})\geq F_{\mathrm{in}}(L^{(k)})+F_{\mathrm{out}}(L^{(k)}).$
By Lemma 5 and Assumption 2, we have that $\frac{|F_{\mathrm{out}}(L)-F_{\mathrm{out}}(L_{\star})|}{F_{\mathrm{in}}(L)-F_{\mathrm{in}}(L_{\star})}\leq 1/3$, and thus
$\left(2+\frac{1}{3}\right)F_{\mathrm{in}}(L^{(l)})\geq\left(1-\frac{1}{3}\right)F_{\mathrm{in}}(L^{(k)}),$
which simplifies to
$5F_{\mathrm{in}}(L^{(l)})\geq F(L^{(k)}).$
Therefore, by applying (40) with $L=L^{(k)}$ and $L^{\prime}=L^{(l)}$, we obtain (41), completing the proof of (41) for case 2.
Finally, observe that $L^{(k+1)}$ is the span of the top $d$ eigenvectors of the matrix
$\sum_{\boldsymbol{x}\in\mathcal{X}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})}=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})}+\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})}.$
Note that the outlier contribution can be bounded as
$\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})}\preceq\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\epsilon_{k}},$
and the inlier contribution satisfies
$\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})}\succeq\sum_{\begin{subarray}{c}\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}\\ \operatorname{dist}(\boldsymbol{x},L^{(k)})\leq\beta\epsilon_{k}\end{subarray}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})}\succeq\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\beta\epsilon_{k}},$
where $\mathcal{X}_{\mathrm{in}}^{\prime}$ is the subset of inliers defined in (41).
Define
$\boldsymbol{A}:=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})},\quad\boldsymbol{B}:=\sum_{\boldsymbol{x}\in\mathcal{X}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\operatorname{dist}(\boldsymbol{x},L^{(k)}),\epsilon_{k})}.$
Then by combining the inequalities above, along with (14) and (41), we obtain
$\sin\theta_{0}\cdot\sigma_{d}(\boldsymbol{A})\geq 2\cdot\sigma_{1}(\boldsymbol{B}-\boldsymbol{A}).$
[…]
$\sin(\theta_{1}(L^{(k+1)},L_{\star}))=\|\boldsymbol{E}\boldsymbol{F}\|\leq\frac{\sigma_{1}(\boldsymbol{B}-\boldsymbol{A})}{\delta}=\frac{2\cdot\sigma_{1}(\boldsymbol{B}-\boldsymbol{A})}{\sigma_{d}(\boldsymbol{A})}\leq\sin\theta_{0}.$
This concludes the proof.
Proof of Lemma 5.
We again denote the principal angles between $L$ and $L_{\star}$ by $\{\theta_{j}\}_{j=1}^{d}$. We first claim that
$F_{\mathrm{in}}(L)$ $\geq\frac{\sum_{j=1}^{d}\sin\theta_{j}}{\sqrt{d}}S_{\mathrm{in}},$ (42)
$|F_{\mathrm{out}}(L)-F_{\mathrm{out}}(L_{\star})|$ $\leq S_{\mathrm{out}}\sum_{j=1}^{d}\theta_{j}.$ (43)
Then the lemma follows from (42) and (43) using the fact that $F_{\mathrm{in}}(L_{\star})=0$.
To prove (42), we again let the principal vectors between $L$ and $L_{\star}$ be $\boldsymbol{u}_{1},\cdots,\boldsymbol{u}_{d}$ and $\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{d}$. Then any $\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}$ can be written as
$\boldsymbol{x}=\sum_{j=1}^{d}x_{j}\boldsymbol{v}_{j},\text{where $x_{j}=\boldsymbol{v}_{j}^{\top}\boldsymbol{x}$},$
and the projection to $L$ can be written as
$P_{L}(\boldsymbol{x})=\sum_{j=1}^{d}x_{j}P_{L}(\boldsymbol{v}_{j}).$
Note that $\boldsymbol{v}_{j}$ has an angle of $\theta_{j}$ with $L$ and $\|\boldsymbol{v}_{j}-P_{L}(\boldsymbol{v}_{j})\|=\sin\theta_{j}$. Consequently,
$\operatorname{dist}(\boldsymbol{x},L)=\|\boldsymbol{x}-\boldsymbol{P}_{L}\boldsymbol{x}\|=\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}x_{j}^{2}}=\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}(\boldsymbol{v}_{j}^{\top}\boldsymbol{x})^{2}}\geq\frac{1}{\sqrt{d}}\sum_{j=1}^{d}\sin\theta_{j}|\boldsymbol{v}_{j}^{\top}\boldsymbol{x}|.$
Summing the above inequality over all $\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}$ and applying the definition of $S_{\mathrm{in}}$, (42) is proved.
To prove (43), let $\{\boldsymbol{u}_{1},\cdots,\boldsymbol{u}_{d},\boldsymbol{w}_{1}\}\in\mathbb{R}^{D}$ be an orthonormal set, and let $L(t)=\mathrm{Sp}(\cos t\theta\boldsymbol{u}_{1}+\sin t\theta\boldsymbol{w}_{1},\boldsymbol{u}_{2},\cdots,\boldsymbol{w}_{d})$. We first notice that
$\left|\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\operatorname{dist}(\boldsymbol{x},L(t))\Big{|}_{t=0}\right|$ $=2\theta\left|\frac{(\cos t\theta\boldsymbol{u}_{1}+\sin t\theta\boldsymbol{w}_{1})^{\top}\boldsymbol{x}\boldsymbol{x}^{\top}(-\sin t\theta\boldsymbol{u}_{1}+\cos t\theta\boldsymbol{w}_{1})}{\operatorname{dist}(\operatorname{dist}(\boldsymbol{x},L(t)))}\right|$ (44)
$\leq 2\theta\left\|\frac{\boldsymbol{P}_{L(t)}\boldsymbol{x}\boldsymbol{x}^{\top}\boldsymbol{P}_{L(t)^{\perp}}}{\operatorname{dist}(\boldsymbol{x},L(t))}\right\|.$
Now let $\boldsymbol{u}_{1},\boldsymbol{u}_{2},\cdots,\boldsymbol{u}_{d}$ and $\boldsymbol{v}_{1},\boldsymbol{v}_{2},\cdots,\boldsymbol{v}_{d}$ be the principal vectors between $L_{\star}$ and $L$ with principal angles $\theta_{1},\cdots,\theta_{d}$, and define that for all $1\leq k\leq d$,
$L_{k}=\mathrm{Span}(\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{k},\boldsymbol{u}_{k+1},\cdots,\boldsymbol{u}_{d}).$
In particular, $L_{0}=L_{\star}$ and $L_{d}=L$. Then $\operatorname{dist}(L_{k-1},L_{k})=\theta_{k}$ and integrating $\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\operatorname{dist}(\boldsymbol{x},L(t))$ and applying (44) implies that
$\left|\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\operatorname{dist}(\boldsymbol{x},L_{k})-\operatorname{dist}(\boldsymbol{x},L_{k-1})\right|\leq\theta_{k}S_{\mathrm{out}}.$
Combining the above inequality for all $0\leq k\leq d-1$, (43) is proved.
Proof of Lemma 6.
We begin by noting that $\frac{F_{\mathrm{in}}(L)}{|\mathcal{X}_{\mathrm{in}}|}=\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\operatorname{dist}(\boldsymbol{x},L)$. Combining this with the fact that $\operatorname{dist}(\boldsymbol{x},L)\leq\sin\theta_{1}(L,L_{\star})\|\boldsymbol{x}\|$ yields
$\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\operatorname{dist}(\boldsymbol{x},L)\leq\sin\theta_{1}(L,L_{\star})\cdot\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\|\boldsymbol{x}\|.$ (45)
To prove the lemma, it is thus sufficient to prove that there is a $c_{3}>0$ such that
$\sup_{L\in\mathscr{G}(D,d)}\frac{|\{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},L)<c_{3}\cdot\sin\theta_{1}(L,L_{\star})\cdot\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\|\boldsymbol{x}\|\}|}{|\mathcal{X}|}<\gamma.$ (46)
This follows from the definition of $q_{\gamma}$ and the fact that (45) implies the following is a sufficient condition for the lemma:
$c_{3}\cdot\sin\theta_{1}(L,L_{\star})\cdot\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\|\boldsymbol{x}\|\leq q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}}).$
To prove the existence of such a $c_{3}$, we will show that Assumption 1 implies that
$\frac{q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}})}{\sin\theta_{1}(L,L_{\star})}$
is uniformly lower bounded by a positive constant for $L\neq L_{\star}$.
We prove this by contradiction. Suppose the statement does not hold. Then there exists a sequence $\{L^{(k)}\}_{k\geq 1}$ such that
$\lim_{k\to\infty}\frac{q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L^{(k)})\}_{\boldsymbol{x}\in\mathcal{X}})}{\sin\theta_{1}(L^{(k)},L_{\star})}=0.$ (47)
Since the Grassmannian is compact, we may assume without loss of generality that $L^{(k)}\to\hat{L}$ (otherwise, we can choose a convergent subsequence of $L^{(k)}$).
Case 1: $\theta_{1}(\hat{L},L_{\star})\neq 0$. Then $\hat{L}\neq L_{*}$ and (47) implies $q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},\hat{L})\}_{\boldsymbol{x}\in\mathcal{X}})=0,$ which contradicts Assumption 1, which claims that fewer than $\gamma$ fraction of points can lie in the $d$-dimensional subspace $\hat{L}$.
Case 2: $\theta_{1}(\hat{L},L_{\star})=0$, i.e., $\hat{L}=L_{\star}$. Without loss of generality, we may assume the principal vector $\boldsymbol{v}_{1}(L^{(k)},L_{\star})$ converges to some $\hat{\boldsymbol{v}}_{1}\in L_{\star}$ (otherwise, choose a subsequence with this property). Then, for any $\boldsymbol{x}\in L_{\star}$,
$\liminf_{k\to\infty}\frac{\operatorname{dist}(\boldsymbol{x},L^{(k)})^{2}}{\sin^{2}\theta_{1}(L^{(k)},L_{\star})}=\liminf_{k\to\infty}\left(|\boldsymbol{x}^{\top}\boldsymbol{v}_{1}(L^{(k)},L_{\star})|^{2}+\sum_{j=2}^{d}\frac{\sin^{2}\theta_{j}(L^{(k)},L_{\star})}{\sin^{2}\theta_{1}(L^{(k)},L_{\star})}|\boldsymbol{x}^{\top}\boldsymbol{v}_{j}(L^{(k)},L_{\star})|^{2}\right)$
$\geq$ $\liminf_{k\to\infty}|\boldsymbol{x}^{\top}\boldsymbol{v}_{1}(L^{(k)},L_{\star})|^{2}=|\boldsymbol{x}^{\top}\hat{\boldsymbol{v}}_{1}|^{2}.$
On the other hand, for $\boldsymbol{x}\notin L_{\star}$, $\operatorname{dist}(\boldsymbol{x},L^{(k)})/\sin\theta_{1}(L^{(k)},L_{\star})\to\infty$ as $k\to\infty$.
Thus, (47) can occur only if at least $\gamma|\mathcal{X}|$ points satisfy $\boldsymbol{x}^{\top}\hat{\boldsymbol{v}}_{1}=0$ and $\boldsymbol{x}\in L_{*}$, i.e., $\gamma|\mathcal{X}|$ points lie in the (d1)(d-1)-dimensional subspace $L_{\star}\cap\mathrm{Sp}(\hat{\boldsymbol{v}}_{1})^{\perp}$, which again contradicts Assumption 1.
In summary, the existence of $c_{3}$ follows by contradiction through the analysis of two cases.
First, note that as $L\to L_{\star}$,
$q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}})\asymp q_{\gamma\frac{|\mathcal{X}|}{|\mathcal{X}_{\mathrm{in}}|}}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}).$
so that in a neighborhood around $L_{\star}$, we only need to understand the behavior of (47) with respect to the inlier points.
Next, observe that for $\boldsymbol{x}\in L_{\star}$, $\frac{\operatorname{dist}(\boldsymbol{x},L)^{2}}{\sin^{2}\theta_{1}(L,L_{\star})}=|\boldsymbol{x}^{\top}\boldsymbol{v}_{j}(L,L_{\star})|^{2}+\sum_{j=2}^{d}\frac{\sin^{2}\theta_{j}(L,L_{\star})}{\sin^{2}(\theta_{1}(L,L_{\star})}|\boldsymbol{x}^{\top}\boldsymbol{v}_{j}(L,L_{\star})|^{2}$, where $\boldsymbol{v}_{j}(L,L_{\star})$ is a principal vector for $L_{\star}$ relative to $L$. Consequently,
$\{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}:$ $\operatorname{dist}(\boldsymbol{x},L)<c_{3}\cdot\sin\theta_{1}(L,L_{\star})\cdot\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\|\boldsymbol{x}\|\}$
$\subseteq\{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}:|\boldsymbol{x}^{\top}\boldsymbol{v}_{1}(L,L_{\star})|<c_{3}\cdot\mathrm{mean}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\|\boldsymbol{x}\|\}.$
By Assumption 1, we thus know that there is a $c_{3}$ such that ${q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}})}/{\sin\theta_{1}(L,L_{\star})}>c_{3}$ for all $L\in B(L_{\star},\delta)$.
Define $L_{0}=\mathrm{Sp}(\boldsymbol{v}_{2},\dots,\boldsymbol{v}_{d})$. Then, by Assumption 1, we know that less than a $\gamma$ fraction of points lie in $L_{0}\cup L$. In particular, for $L\neq L_{\star}$,
$\frac{q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}})}{\sin\theta_{1}(L,L_{\star})}>\min\left(\left\{|\boldsymbol{v}_{1}^{\top}\boldsymbol{x}|:\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}\setminus L_{0}\right\}\cup\left\{\frac{\operatorname{dist}(\boldsymbol{x},L)}{\sin\theta_{1}(L,L_{\star})}:\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}\setminus L\right\}\right)$
Since $\frac{q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}})}{\sin\theta_{1}(L,L_{\star})}\not\to 0$ as $L\to L_{\star}$ and $\frac{q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}})}{\sin\theta_{1}(L,L_{\star})}$, there exists a $c_{3}$ such that the lemma holds.
6.2 Proof of Theorem 2
We begin with some background before proceeding to the proof of the theorem.
6.2.1 Background of proof
A parametrization of paths in $\mathscr{A}(D,d)$. Recall that the underlying affine subspace in our model is $A_{\star}]=[(L_{\star},0)]$, where we choose the representative $(L_{\star},0)$ WLOG.
For another affine subspace parametrized by $A=(L,\boldsymbol{m})$, let us consider a path between $A_{\star}$ and $A$ defined as follows: let the principal vectors of $L_{\star}$ and $L$ be
$L_{\star}=\mathrm{Sp}(\boldsymbol{u}_{1},\cdots,\boldsymbol{u}_{d}),\,\,L=\mathrm{Sp}(\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{d}),$
where $\boldsymbol{v}_{i}=\cos\theta_{i}\boldsymbol{u}_{i}+\sin\theta_{i}\boldsymbol{w}_{i}$ such that $\boldsymbol{u}_{i}\perp\boldsymbol{w}_{j}$ for all $1\leq i,j\leq d$.
Then a path of affine subspaces from $A_{\star}$ to $A$ given by $C:[0,1]\rightarrow\mathscr{A}(D,d)$ is defined by
$C(t)=(1-t)\boldsymbol{a}_{\star}+t\boldsymbol{a}+\mathrm{Sp}(\{\cos t\theta_{i}\boldsymbol{u}_{i}+\sin t\theta_{i}\boldsymbol{w}_{i}\}_{i=1}^{d}),$ (48)
where $\boldsymbol{a}_{\star}$ is the point in $A_{\star}$ closest to $A$ and $\boldsymbol{a}$ is the point on $A$ closest to $A_{\star}$. Note that $\boldsymbol{a}-\boldsymbol{a}_{\star}\perp\mathrm{Sp}(L\oplus L_{\star})$. Denote $\boldsymbol{b}=\boldsymbol{a}-\boldsymbol{a}_{\star}$, then
$C(t)=\boldsymbol{a}_{\star}+t\boldsymbol{b}+\mathrm{Sp}(\{\cos t\theta_{i}\boldsymbol{u}_{i}+\sin t\theta_{i}\boldsymbol{w}_{i}\}_{i=1}^{d}).$ (49)
Note that here
$\operatorname{dist}^{2}(0,A)$ $=\operatorname{dist}^{2}(-\boldsymbol{a},L)=\operatorname{dist}^{2}(-\boldsymbol{a}_{\star}-\boldsymbol{b},L)$ (50)
$=$ $\|\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}\operatorname{dist}^{2}(-\boldsymbol{P}_{\mathrm{Sp}(\boldsymbol{u}_{i},\boldsymbol{w}_{i})}\boldsymbol{a}_{\star},\mathrm{Sp}(\{\cos\theta_{i}\boldsymbol{u}_{i}+\sin\theta_{i}\boldsymbol{w}_{i}\}_{i=1}^{d})$ (51)
${=}\|\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}(\boldsymbol{a}_{\star}^{\top}\boldsymbol{u}_{i})^{2}\sin^{2}\theta_{i},$
where the last step applies that $\boldsymbol{a}_{\star}\perp\boldsymbol{w}_{i}$. Therefore the definition in Section 4.2 gives
$\operatorname{dist}^{2}(A_{\star},[A])$ $=\operatorname{dist}^{2}(0,[A])+\sum_{j}\theta_{j}^{2}$ (52)
$=\|\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}\big{(}(\boldsymbol{a}_{\star}^{\top}\boldsymbol{u}_{i})^{2}\sin^{2}\theta_{i}+\theta_{i}^{2}\big{)}.$
Auxiliary functions. Similar to the proof of Theorem 1, the proof of Theorem 2 also depends on $F_{\epsilon}^{(a)}(A)$, a regularized objective function $F^{(a)}$ defined in (9):
$F_{\epsilon}^{(a)}(A)$ $=\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},A)>\epsilon}\operatorname{dist}(\boldsymbol{x},A)+\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},A)\leq\epsilon}\left(\frac{1}{2}\epsilon+\frac{\operatorname{dist}(\boldsymbol{x},A)^{2}}{2\epsilon}\right).$ (53)
Define
$G_{\epsilon_{k}}^{(a)}(A,A_{0})$ $=\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},A_{0})>\epsilon_{k}}\left(\frac{1}{2}\operatorname{dist}(\boldsymbol{x}_{i},A_{0})+\frac{\operatorname{dist}(\boldsymbol{x}_{i},A)^{2}}{2\operatorname{dist}(\boldsymbol{x}_{i},A_{0})}\right)+$ (54)
$\sum_{\boldsymbol{x}\in\mathcal{X}:\operatorname{dist}(\boldsymbol{x},A_{0})\leq\epsilon_{k}}\left(\frac{1}{2}\epsilon_{k}+\frac{\operatorname{dist}(\boldsymbol{x}_{i},A)^{2}}{2\epsilon}\right).$
6.2.2 Proof of Theorem 2
We prove that the algorithm ensures the sequence $F^{(a)}_{\epsilon_{k}}(A^{(k)})$ is nonincreasing over iterations. That is, we aim to generalize Lemma 2 to the affine setting.
Lemma 8 (Decrease over iterations, affine case).
$F_{\epsilon}(A)-F_{\epsilon}(T^{(a)}(A))\geq c\epsilon\operatorname{dist}(A_{\star},[A]).$ (55)
Proof of Lemma 8.
Let us first investigate $\operatorname{dist}^{2}(\boldsymbol{x},C(t))$. Since $C(t)$ defined in (48) can be decomposed into the direct sum of $d+1$ components (note that by definition, $\boldsymbol{P}_{\mathrm{Sp}(L\oplus L_{\star})}\boldsymbol{a}=\boldsymbol{P}_{\mathrm{Sp}(L\oplus L_{\star})}\boldsymbol{a}_{\star}$):
$\boldsymbol{P}_{\mathrm{Sp}(\boldsymbol{u}_{i},\boldsymbol{w}_{i})}C(t)$ $=(\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}+\boldsymbol{w}_{i}\boldsymbol{w}_{i}^{\top})\boldsymbol{a}_{\star}+\mathrm{Sp}(\cos t\theta_{i}\boldsymbol{u}_{i}+\sin t\theta_{i}\boldsymbol{w}_{i}),1\leq i\leq d$ (56)
$\boldsymbol{P}_{\mathrm{Sp}(L\oplus L_{\star})^{\perp}}C(t)$ $=t\boldsymbol{b},$ (57)
the distances $\operatorname{dist}^{2}(\boldsymbol{x},C(t))$ can also be decomposed using
$\operatorname{dist}^{2}(\boldsymbol{x},C(t))$
$=\|\boldsymbol{P}_{\mathrm{Sp}(L\oplus L_{\star})^{\perp}}\boldsymbol{x}-t\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}\operatorname{dist}^{2}(\boldsymbol{P}_{\mathrm{Sp}(\boldsymbol{u}_{i},\boldsymbol{w}_{i})}\boldsymbol{x},\boldsymbol{P}_{\mathrm{Sp}(\boldsymbol{u}_{i},\boldsymbol{w}_{i})}C(t))$
$=\|\boldsymbol{P}_{\mathrm{Sp}(L\oplus L_{\star})^{\perp}}\boldsymbol{x}-t\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}((\boldsymbol{u}_{i}\sin t\theta_{i}-\boldsymbol{w}_{i}\cos t\theta_{i})^{\top}(\boldsymbol{x}-(\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}+\boldsymbol{w}_{i}\boldsymbol{w}_{i}^{\top})\boldsymbol{a}_{\star}))^{2}$
$=\|\boldsymbol{P}_{\mathrm{Sp}(L\oplus L_{\star})^{\perp}}\boldsymbol{x}-t\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}((\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\sin t\theta_{i}-\boldsymbol{x}^{\top}\boldsymbol{w}_{i}\cos t\theta_{i})^{2},$
where the last step follows from $\boldsymbol{w}_{i}\perp L_{\star}$ and $\boldsymbol{a}_{\star}\in L_{\star}$. Note also that $(\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\sin t\theta_{i}-\boldsymbol{x}^{\top}\boldsymbol{w}_{i}\cos t\theta_{i}=\sin(t\theta_{i}-\theta^{\prime}_{i})\sqrt{((\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i})^{2}+(\boldsymbol{x}^{\top}\boldsymbol{w}_{i})^{2}}$ for some $\theta^{\prime}_{i}$
We proceed by studying $\varphi(t)=G_{\epsilon}^{(a)}(C(t),A)$. In particular, we will bound the first and second derivatives of $\varphi$. Since the second derivative of $\sin^{2}x$ is bounded in [2,2][-2,2], we have
$\frac{{\,\mathrm{d}}^{2}}{{\,\mathrm{d}}t^{2}}\operatorname{dist}^{2}(\boldsymbol{x},C(t))$ $\leq 2\|\boldsymbol{b}\|^{2}+2\sum_{i=1}^{d}\theta_{i}^{2}\Big{[}((\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i})^{2}+(\boldsymbol{x}^{\top}\boldsymbol{w}_{i})^{2}\Big{]}$
$\leq 2\|\boldsymbol{b}\|^{2}+2\theta_{1}^{2}\|\boldsymbol{x}-\boldsymbol{a}_{\star}\|^{2}\leq 2\|\boldsymbol{b}\|^{2}+4\theta_{1}^{2}(\|\boldsymbol{x}\|^{2}+\|\boldsymbol{a}_{\star}\|^{2})\leq.$...
$\leq 2\|\boldsymbol{b}\|^{2}+2\sum_{i=1}^{d}\theta_{i}^{2}\left[2(\boldsymbol{a}_{\star}^{\top}\boldsymbol{u}_{i})^{2}+2(\boldsymbol{x}^{\top}\boldsymbol{u}_{i})^{2}+(\boldsymbol{x}^{\top}\boldsymbol{w}_{i})^{2}\right]$
$\leq 4\|\boldsymbol{x}\|^{2}\sum_{i=1}^{n}\theta_{i}^{2}+10\operatorname{dist}^{2}(0,[A])$
$=10\max(\|\boldsymbol{x}\|^{2},1)\operatorname{dist}^{2}(A_{\star},[A]),$
where the last two steps apply (50) and $\theta\leq\pi\sin\theta/2$.
Combining the above estimation for all $\boldsymbol{x}\in\mathcal{X}$, we have
$\varphi^{\prime\prime}(t)=\frac{{\,\mathrm{d}}^{2}}{{\,\mathrm{d}}t^{2}}G^{(a)}_{\epsilon}(C(t),A)\leq\frac{C}{\epsilon}\operatorname{dist}^{2}(A_{\star},[A])$
for some $C$ depending on $\mathcal{X}$.
As for the first derivative of $\operatorname{dist}^{2}(\boldsymbol{x},C(t))$ at $t=1$, we have
$\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\operatorname{dist}^{2}(\boldsymbol{x},C(t))\Big{|}_{t=1}=\boldsymbol{b}^{\top}(\boldsymbol{P}_{\mathrm{Sp}(L\oplus L_{\star})^{\perp}}\boldsymbol{x}-\boldsymbol{b})$
$+\sum_{i=1}^{d}\theta_{i}\Big{(}(\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\sin\theta_{i}-\boldsymbol{x}^{\top}\boldsymbol{w}_{i}\cos\theta_{i}\Big{)}\Big{(}(\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\cos\theta_{i}+\boldsymbol{x}^{\top}\boldsymbol{w}_{i}\sin\theta_{i}\Big{)}.$
Combining it with
$\operatorname{dist}(\boldsymbol{x},C(1))=\sqrt{\Big{\|}\boldsymbol{P}_{\mathrm{Sp}(L\oplus L_{\star})^{\perp}}\boldsymbol{x}-\boldsymbol{b}\Big{\|}^{2}+\sum_{i=1}^{d}\Big{(}(\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\sin\theta_{i}-\boldsymbol{x}^{\top}\boldsymbol{w}_{i}\cos\theta_{i}\Big{)}^{2}},$
we have
$\frac{\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\operatorname{dist}^{2}(\boldsymbol{x},C(t))\Big{|}_{t=1}}{\operatorname{dist}(\boldsymbol{x},C(1))}\leq$ $\sqrt{\|\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}\theta_{i}^{2}\Big{(}(\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\cos\theta_{i}+\boldsymbol{x}^{\top}\boldsymbol{w}_{i}\sin\theta_{i}\Big{)}^{2}}$
$\leq$ $\sqrt{\|\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}\theta_{i}^{2}\Big{(}(\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\Big{)}^{2}}+\sqrt{\sum_{i=1}^{d}\theta_{i}^{4}(\boldsymbol{x}^{\top}\boldsymbol{w}_{i})^{2}}$
$\leq$ $\frac{\pi}{2}\operatorname{dist}(\boldsymbol{P}_{A_{\star}}\boldsymbol{x},[A])+\sum_{i=1}^{d}\theta_{i}^{2}(\boldsymbol{x}^{\top}\boldsymbol{w}_{i})^{2}.$
On the other hand, if we restrict that $\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}$, then
$\operatorname{dist}^{2}(\boldsymbol{x},C(1))$ $=\operatorname{dist}^{2}(\boldsymbol{x},[A])=\|\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}\Big{(}(\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\Big{)}^{2}\sin^{2}\theta_{i}$
$\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\operatorname{dist}^{2}(\boldsymbol{x},C(t))\Big{|}_{t=1}$ $=\|\boldsymbol{b}\|^{2}+\sum_{i=1}^{d}\theta_{i}\Big{(}(\boldsymbol{x}-\boldsymbol{a}_{\star})^{\top}\boldsymbol{u}_{i}\Big{)}^{2}\sin\theta_{i}\cos\theta_{i}\geq\cos^{2}\theta_{1}\operatorname{dist}^{2}(\boldsymbol{x},C(1)).$
Combining the estimations above, along with
$\sum_{\boldsymbol{x}\in\mathcal{X}}\sum_{i=1}^{d}\theta_{i}^{2}(\boldsymbol{x}^{\top}\boldsymbol{w}_{i})^{2}\leq\sum_{i=1}^{d}\theta_{i}^{2}\|\boldsymbol{P}_{L_{\star}^{\perp}}\boldsymbol{X}_{\mathrm{out}}\|^{2},$
we have that for $\varphi(t)=G_{\epsilon}^{(a)}(C(t),A)$,
$\varphi^{\prime}(1)\geq\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\frac{\cos^{2}\theta_{1}\operatorname{dist}^{2}(\boldsymbol{x},A)}{2\max(\operatorname{dist}(\boldsymbol{x},A),\epsilon)}-\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\Big{(}\frac{\pi}{2}\operatorname{dist}(\boldsymbol{P}_{A_{\star}}\boldsymbol{x},[A])+\theta_{1}^{2}\|\boldsymbol{P}_{L_{\star}^{\perp}}\boldsymbol{x}\|^{2}\Big{)}.$
Recall from definition of $F,G$, we have
$\min_{t}\varphi(t)\geq G_{\epsilon}(T^{(a)}(A),A)\geq F_{\epsilon}(T^{(a)}(A)),$
and combining it with Assumption 5 and the estimation of $\varphi^{\prime\prime}(t)$ and $\varphi^{\prime}(1)$ above, we have
$F_{\epsilon}(A)-F_{\epsilon}(T^{(a)}(A))\geq\varphi(1)-\min_{t}\varphi(t)\geq\frac{(\varphi^{\prime}(1))^{2}}{2\max_{t}\varphi^{\prime\prime}(t)}\geq c\epsilon\operatorname{dist}(A_{\star},[A]),$
and the Lemma is proved. ∎
The following lemma generalizes Lemma 6 and the proof is similar.
Lemma 9.
Under Assumption 4, there exists $c_{3}>0$ such that for all $A]\in\mathscr{A}(D,d)$,
$q_{\gamma}(\{\operatorname{dist}(\boldsymbol{x},[A])\}_{\boldsymbol{x}\in\mathcal{X}})\geq c_{3}\frac{F^{(a)}_{\mathrm{in}}(A)}{|\mathcal{X}_{\mathrm{in}}|}.$ (58)
The next lemma shows that $\operatorname{dist}(A_{\star},[A])$ serves as a lower bound for the FMS objective over at least half of the inliers:
Lemma 10.
Under Assumption 3, for any subset $\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}$ with $|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\frac{1}{2}|\mathcal{X}_{\mathrm{in}}|$, we have
$\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\operatorname{dist}(\boldsymbol{x},A)\geq c\cdot\operatorname{dist}(A_{\star},[A]),$ (59)
for some constant $c>0$.
Proof of Lemma 10.
Let $\boldsymbol{b}$ denote the projection of the origin onto the affine subspace $A_{\star}$. For each inlier $\boldsymbol{x}\in A_{\star}$, we can express
$\boldsymbol{x}=\sum_{j=1}^{d}a_{j}\boldsymbol{v}_{j}+\boldsymbol{b},$
where $\{\boldsymbol{v}_{j}\}_{j=1}^{d}$ form an orthonormal basis for the linear part of $A_{\star}$. Let $\{\boldsymbol{u}_{j}\}_{j=1}^{d}$ and $\{\boldsymbol{w}_{j}\}_{j=1}^{d}$ be orthonormal vectors such that each $\boldsymbol{u}_{j}$ can be decomposed as
$\boldsymbol{u}_{j}=\cos\theta_{j}\boldsymbol{v}_{j}+\sin\theta_{j}\boldsymbol{w}_{j},$
with $\boldsymbol{w}_{j}\perp A_{\star}$. Then the squared distance from $\boldsymbol{x}$ to $A$ can be written as
$\operatorname{dist}(\boldsymbol{x},A)^{2}=\sum_{j=1}^{d}\left(\sin\theta_{j}\cdot\boldsymbol{x}^{\top}\boldsymbol{u}_{j}+\boldsymbol{b}^{\top}\boldsymbol{w}_{j}\right)^{2}.$
Note that the distance from $A_{\star}$ to $[A]$ satisfies
$\operatorname{dist}(A_{\star},[A])^{2}=\|\boldsymbol{b}\|^{2}+\sum_{j=1}^{d}\theta_{j}^{2}.$
Now suppose, for contradiction, that the claim does not hold. Then there exists an affine subspace $A\neq A_{\star}$ such that $\operatorname{dist}(A_{\star},[A])\neq 0$, yet
$\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\operatorname{dist}(\boldsymbol{x},A)=0.$
This implies that all points in $\mathcal{X}_{\mathrm{in}}^{\prime}$ lie entirely within $A$, contradicting Assumption 3, which asserts that any affine subspace distinct from $A_{\star}$ cannot contain more than half of the inliers.
Therefore, the bound (59) must hold for some constant $c>0$, completing the proof. ∎
Proof of Theorem 2.
The proof of Theorem 2 follows similarly from the proof of Theorem 1. In particular, Lemma 8 implies that
$F_{\epsilon_{k}}^{(a)}(A^{(k)})-F_{\epsilon_{k}}^{(a)}(T^{(a)}(A^{(k+1)}))\geq C\epsilon_{k}.$
On the other hand, use the same argument as in Lemma 5(a), we have
$F^{(a)}_{in,\epsilon_{k}}(A^{(k)})\leq 2F^{(a)}_{\mathrm{in}}(A_{\star})$
and the argument in Lemma 8 implies that
$|F^{(a)}_{\mathrm{out},\epsilon_{k}}(A^{(k)})-F^{(a)}_{\mathrm{out},\epsilon_{k}}(A_{\star})|\leq F^{(a)}_{\mathrm{in}}(A^{(k)})/2.$
In summary,
$F^{(a)}_{\epsilon_{k}}(A^{(k)})-F^{(a)}(A_{\star})\leq 3F^{(a)}_{\mathrm{in}}(A^{(k)}).$
Combining it with
$F^{(a)}_{\mathrm{in}}(A^{(k)})$ $=\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\operatorname{dist}(\boldsymbol{x},A^{(k)})$
$\leq\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\operatorname{dist}(\boldsymbol{x},L^{(k)})+\operatorname{dist}(A_{\star},[A^{(k)}])$
$\leq\operatorname{dist}(A_{\star},[A^{(k)}])\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}(\|\boldsymbol{x}\|+1)$
and Lemma 9, the same argument as in Theorem 1 can be applied to prove the convergence. ∎
6.3 Proof of Proposition 1
We verify Assumptions 2 and 3 hold with probability approaching 1 in Sections 6.3.1 and 6.3.2, respectively
6.3.1 Verification of Assumption 2
The following lemma gives bounds on the statistics $S_{\mathrm{in}}$ and $S_{\mathrm{out}}$ defined in (10) under this model with high probability. Clearly, given the assumption of the proposition, this lemma implies Assumption 2 holds with fixed $\theta_{0}$, with probability tending to 1 as $n,D\to\infty$.
Lemma 11 (Estimation of parameters under the Haystack model).
Assuming that $n=O(D^{3}\log D)$, $\epsilon\leq 1$, and $d\leq D-2$, then we have the following estimations of parameters in Theorems 1 and Assumption 2 with high probability:
$S_{\mathrm{in}}\gtrsim n_{\mathrm{in}}\sigma_{\mathrm{in}}/\sqrt{d},\,\,\,S_{\mathrm{out}}\lesssim\frac{n_{\mathrm{out}}\sigma_{\mathrm{out}}}{\sqrt{D(D-d)}}.$
Proof of Lemma 11.
For $S_{\mathrm{in}}$, it follows from (Lerman et al., 2015, Lemma 8.2) that
$\Pr\left(S_{\mathrm{in}}\geq\left(\sqrt{\frac{\pi}{2}}n_{\mathrm{in}}-2\sqrt{n_{\mathrm{in}}d}-t\sqrt{n_{\mathrm{in}}}\right)\frac{\sigma_{\mathrm{in}}}{d}\right)\geq 1-e^{-t^{2}/2}$
On the other hand we can use (Maunu et al., 2019, Proposition 12) to show that
$\Pr\left(\sigma_{d}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\frac{\boldsymbol{x}\boldsymbol{x}^{\top}}{\max(\|\boldsymbol{x}\|,\epsilon)}\right)\geq\frac{1}{c}\sqrt{\frac{\sigma_{\mathrm{in}}}{d}}\frac{(1-a)^{2}N_{\text{in}}}{d}+{O}\left(\sqrt{\frac{N_{\text{in}}}{d}}\right)\right)\geq 1-4e^{-c_{1}a^{2}N_{\text{in}}}.$
[…]
Estimation for fixed $L\in G(D,d)$: Assuming that $|\mathcal{X}_{\mathrm{out}}|=n_{\mathrm{out}}$, then Lerman et al. (2015, Proposition 8.3) proves that with probability at least $1-e^{t_{1}^{2}/2}$,
$\|\boldsymbol{P}_{L}\boldsymbol{X}\|\gtrsim\sigma_{\mathrm{out}}(\sqrt{d}+\sqrt{n_{\mathrm{out}}}+t_{1})/\sqrt{D},$
and Lerman et al. (2015, Lemma 8.4) shows that for $\boldsymbol{S}$, a $n_{\mathrm{out}}\times D$ matrix with each row given by $\boldsymbol{P}_{L^{\perp}}\boldsymbol{x}/\|\operatorname{dist}(\boldsymbol{x},L)\|$, with probability at least $1-1.5e^{t_{2}^{2}/2}$
$\|\boldsymbol{S}\|\gtrsim\frac{\sqrt{D-d}+\sqrt{n_{\mathrm{out}}}+t_{2}}{\sqrt{D-d-0.5}}.$
As a result, with probability at least $1-e^{t_{1}^{2}/2}-1.5e^{t_{2}^{2}/2}$ we have
[…]
which is with probability $1-2.5e^{-n_{\mathrm{out}}/2}$ in the order of $\frac{n_{\mathrm{out}}\sigma_{\mathrm{out}}}{\sqrt{D(D-d)}}$ for $n_{\mathrm{out}}\geq D$.
Covering argument: And for any $L_{1},L_{2}\in\mathscr{G}(D,d)$, applying
$\left\|\frac{\boldsymbol{P}_{L_{1}^{\perp}}\boldsymbol{x}}{\operatorname{dist}(\boldsymbol{x},L_{1})}-\frac{\boldsymbol{P}_{L_{2}^{\perp}}\boldsymbol{x}}{\operatorname{dist}(\boldsymbol{x},L_{2})}\right\|<2\min\left(\frac{\|\boldsymbol{P}_{L_{1}^{\perp}}-\boldsymbol{P}_{L_{2}^{\perp}}\|\|\boldsymbol{x}\|}{\operatorname{dist}(\boldsymbol{x},L_{1})},1\right),$
we have
[…]
When $d\leq D-2$, then $2\|\boldsymbol{x}\|\min\Big{(}\frac{\|\boldsymbol{P}_{L_{1}^{\perp}}-\boldsymbol{P}_{L_{2}^{\perp}}\|\|\boldsymbol{x}\|}{\operatorname{dist}(\boldsymbol{x},L_{1})},1\Big{)}$ has expectation bounded by $C\sigma_{\mathrm{out}}\|\boldsymbol{P}_{L_{1}^{\perp}}-\boldsymbol{P}_{L_{2}^{\perp}}\|\sqrt{D/(D-d)}$. So Hoeffding’s inequality in Theorem 2.6.2 of Vershynin (2018) implies that the RHS of (61) is close to its expectation and with high probability that, it is bounded above by
$Cn_{\mathrm{out}}\sigma_{\mathrm{out}}\|\boldsymbol{P}_{L_{1}^{\perp}}-\boldsymbol{P}_{L_{2}^{\perp}}\|\sqrt{D/(D-d)}.$
Combining it with an $\epsilon$-net covering argument with $\epsilon=1/D$ over $L\in\mathscr{G}(D,d)$, the estimation of $S_{\mathrm{out}}$ for a fixed $L$ in (60), and $n_{\mathrm{out}}\approx\alpha_{0}n$, the estimation of $S_{\mathrm{out}}$ is in the order of ${n_{\mathrm{out}}\sigma_{\mathrm{out}}}/{\sqrt{D(D-d)}}$ with probability $1-CD^{D(D-d)}e^{-n_{\mathrm{out}}/2}$, which proves Lemma 11. ∎
6.3.2 Verification of Assumption 3 under the Generalized Haystack Model
We begin by presenting a lemma that introduces and reviews several key probability bounds for generalized chi-squared and normal distributions, which will be used in the verification.
Lemma 12.
(a) Assuming that $\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n}\sim N(0,\boldsymbol{\Sigma})$ with $\boldsymbol{\Sigma}\in\mathbb{R}^{p\times p}$, then
$\Pr\left(\frac{1}{n}\sum_{i=1}^{n}\|\boldsymbol{x}_{i}\|\leq\sqrt{2\mathrm{tr}(\boldsymbol{\Sigma})}\right)\geq 1-e^{-\frac{n\mathrm{tr}(\boldsymbol{\Sigma})}{8}}.$
(b) Assuming that $x_{1},\cdots,x_{n}$ are $n$ samples of a random variable $X$, then its empirical quantile are not far away from its population quantile:
$\Pr\left(q_{\alpha_{0}}(\{x_{i}\}_{i=1}^{n})\geq q_{\alpha_{0}/2}(X)\right)\geq 1-e^{-n\alpha_{0}^{2}/2},\,\,\Pr\left(q_{1-\alpha_{0}}(\{\|x_{i}\|\}_{i=1}^{n})\leq q_{1-\alpha_{0}/2}(X)\right)\geq 1-e^{-n\alpha_{0}^{2}/2}.$
(c) Assuming that $\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n}\sim N(0,\boldsymbol{\Sigma})$ with $\boldsymbol{\Sigma}\in\mathbb{R}^{p\times p}$, then
$\Pr\left(q_{\alpha_{0}/2}(\{\|\boldsymbol{x}_{i}\|\}_{i=1}^{n})\geq\frac{\alpha_{0}}{4e}\sqrt{2\mathrm{tr}(\boldsymbol{\Sigma})}\right)\geq 1-e^{-n\alpha_{0}^{2}/8}$
(d) Assuming that $\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n}\sim N(0,\boldsymbol{I}_{p\times p})$, then
$\Pr\left(q_{1-\alpha_{0}/2}(\{\|\boldsymbol{x}_{i}\|\}_{i=1}^{n})\leq\sqrt{p}-2\log\frac{\alpha_{0}}{4}\right)\geq 1-e^{-n\alpha_{0}^{2}/8}$
and when $\alpha_{0}/2\leq 1-4\exp(-4/p)$,
$\Pr\left(q_{\alpha_{0}/2}(\{\|\boldsymbol{x}_{i}\|\}_{i=1}^{n})\geq\sqrt{\left(1-\sqrt{\frac{4}{p}\log\frac{4}{\alpha_{0}}}\right)p}\right)\geq 1-e^{-n\alpha_{0}^{2}/8}$
(e) If $g_{1},\cdots,g_{n}$ are from $N(0,1)$, then for $0\leq\alpha_{0}\leq 1$,
$\Pr\left(q_{\alpha_{0}}(\{|g_{i}|\}_{i=1}^{n})\geq\alpha_{0}/2\right)\geq 1-e^{-n\alpha_{0}^{2}/2}$
and
$\Pr\left(\sum_{i=1}^{n}g_{i}^{2}\cdot I(g_{i}\leq\alpha_{0}/2)\geq\frac{1}{25}n\alpha_{0}^{3}\right)\geq 1-e^{-2n\alpha_{0}^{2}/625}.$
We will first assume Lemma 12 is true and verify (12), (13) and (14). We then prove the lemma.
Verification of (12): Let $\boldsymbol{z}=(\boldsymbol{U}_{L_{\star}}^{\top}\boldsymbol{\Sigma}_{in}\boldsymbol{U}_{L_{\star}})^{-1/2}\boldsymbol{U}_{L_{\star}}^{\top}\boldsymbol{x}$ and $\boldsymbol{W}=\boldsymbol{U}_{L}^{\top}\boldsymbol{U}_{L_{\star}}$, then $\operatorname{dist}(\boldsymbol{x},L)=\|\boldsymbol{P}_{L}\boldsymbol{x}\|=\|\boldsymbol{W}\boldsymbol{z}\|$ and $\boldsymbol{z}\sim\mathcal{N}(0,\boldsymbol{I}_{d\times d})$. As a result, it is equivalent to prove that for $\boldsymbol{z}_{1},\boldsymbol{z}_{n_{\mathrm{in}}}\sim\mathcal{N}(0,\boldsymbol{I}_{d\times d})$,
$\beta_{1}\cdot\mathrm{mean}(\{\|\boldsymbol{W}\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})\leq q_{\alpha_{0}}(\{\|\boldsymbol{W}\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})$ (62)
holds for all $\boldsymbol{W}\in\mathbb{R}^{d\times d}$ with $\|\boldsymbol{W}\|_{F}=1$.
It will be proved by an $\epsilon$-net argument over all such $\boldsymbol{W}$. For any fixed $\boldsymbol{W}$, by Lemma 12(a)(c) (let $\boldsymbol{\Sigma}=\boldsymbol{W}^{\top}\boldsymbol{W}$ and therefore $\mathrm{tr}(\boldsymbol{\Sigma})=1$), we have
$\Pr\left(\frac{8e}{\alpha_{0}}q_{\alpha_{0}/2}(\{\|\boldsymbol{W}\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})-\mathrm{mean}(\{\|\boldsymbol{W}\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})\geq\sqrt{2}\right)\geq 1-e^{-n_{\mathrm{in}}\alpha_{0}^{2}/8}-e^{-n_{\mathrm{in}}/8}.$
Then apply the $\epsilon$-net argument to set of all $d\times d$ matrices $\boldsymbol{W}$ using the perturbation argument:
$\mathrm{mean}(\{\|\boldsymbol{W}^{\prime}\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})\leq\mathrm{mean}(\{\|\boldsymbol{W}\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})+\|\boldsymbol{W}-\boldsymbol{W}^{\prime}\|\mathrm{mean}(\{\|\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})$ (63)
and by the property of quantile,
$q_{\alpha_{0}}(\{\|\boldsymbol{W}^{\prime}\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})\geq q_{\alpha_{0}/2}(\{\|\boldsymbol{W}\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})-\|\boldsymbol{W}-\boldsymbol{W}^{\prime}\|q_{1-\alpha_{0}/2}(\{\|\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}}).$ (64)
Also, applying Lemma 12(a)(c), we have that with probability $1-e^{n_{\mathrm{in}}/8}-e^{n_{\mathrm{in}}\alpha_{0}^{2}/8}$,
$\mathrm{mean}(\{\|\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})$ $\leq\sqrt{2d}$ (65)
$q_{1-\alpha_{0}/2}(\{\|\boldsymbol{z}_{i}\|\}_{1\leq i\leq n_{\mathrm{in}}})$ $\leq\sqrt{d}-2\log\frac{\alpha_{0}}{4}$ (66)
Use an $\epsilon$-net argument with $\epsilon=\frac{\sqrt{2}}{\frac{8e}{\alpha_{0}}\sqrt{2d}+\sqrt{d}-2\log\frac{\alpha_{0}}{4}}$, we proved that for $\beta_{1}=8e/\alpha_{0}$, (12) holds with probability $1-(\frac{\sqrt{2}}{\frac{8e}{\alpha_{0}}\sqrt{2d}+\sqrt{d}-2\log\frac{\alpha_{0}}{4}})^{d^{2}}(e^{-n_{\mathrm{in}}/8}+e^{-n_{\mathrm{in}}\alpha_{0}^{2}/8})$.
Verification of (13): For arbitrary PSD matrices $\boldsymbol{A}$, $\boldsymbol{B}$, if $\sigma_{1}(\boldsymbol{A})\leq 1\leq\sigma_{d}(\boldsymbol{B})$, then $\operatorname{dist}(\boldsymbol{A}\boldsymbol{x},L)\leq\operatorname{dist}(\boldsymbol{x},L)\leq\operatorname{dist}(\boldsymbol{B}\boldsymbol{x},L)$. As a result, WLOG it is sufficient to assume that the inliers are sampled from $\mathcal{N}(0,C\sigma_{\text{in}}^{2}\boldsymbol{P}_{L_{\star}}/d)$ and the outliers are sampled from $\mathcal{N}(0,c\sigma_{\text{out}}^{2}\boldsymbol{I}/D)$.
The RHS of (13) can be controlled by Lemma 12(c): with probability $1-\exp(-n_{\mathrm{in}}\alpha_{0}^{2}/2)$,
$q_{\alpha_{0}}\left(\left\{\|\boldsymbol{x}\|\right\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\right)\lesssim{\sigma_{\mathrm{in}}}\frac{\alpha_{0}\sqrt{2}}{2e}.$
As for the LHS of (13), similar to (64), we will use the perturbation bound
$q_{\frac{\gamma-\alpha_{0}\gamma_{\star}}{1-\gamma_{\star}}}(\{\operatorname{dist}(\boldsymbol{x},L^{\prime})\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}})\geq q_{\frac{\gamma-\alpha_{0}\gamma_{\star}}{2(1-\gamma_{\star})}}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}})-\|\boldsymbol{P}_{L}-\boldsymbol{P}_{L^{\prime}}\|q_{1-\frac{\gamma-\alpha_{0}\gamma_{\star}}{2(1-\gamma_{\star})}}(\{\|\boldsymbol{x}\|\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}),$ (67)
and the estimations that for any fixed $L$, Lemma 12(d) implies that with high probability $1-\exp(-n_{\mathrm{out}}\alpha_{1}^{2}/8)$ where $\alpha_{1}=\frac{\gamma-\alpha_{0}\gamma_{\star}}{(1-\gamma_{\star})}$ (note $\alpha_{1}\geq\gamma\geq\min(C/(D-d),1/100)$ and therefore satisfies the condition in Lemma 12(d) for large p=Ddp=D-d) implies
$q_{\frac{\gamma-\alpha_{0}\gamma_{\star}}{2(1-\gamma_{\star})}}(\{\operatorname{dist}(\boldsymbol{x},L)\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}})\gtrsim{\sigma_{\mathrm{out}}}\sqrt{\left(1-\sqrt{\frac{4}{D-d}\log\frac{4}{\alpha_{1}}}\right)\frac{D-d}{D}}\gtrsim\sqrt{\sigma_{\mathrm{out}}\frac{D-d}{D}},$ (68)
and the above still holds for small DdD-d when $\alpha_{1}\geq 1/100$; as well as the estimation also from Lemma 12(d) that
$q_{1-\frac{\gamma-\alpha_{0}\gamma_{\star}}{2(1-\gamma_{\star})}}(\{\|\boldsymbol{x}\|\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}})\leq\frac{{\sigma_{\mathrm{out}}}}{\sqrt{D}}\left(\sqrt{D}-2\log\frac{\gamma-\alpha_{0}\gamma_{\star}}{4(1-\gamma_{\star})}\right)\lesssim{{\sigma_{\mathrm{out}}}}.$ (69)
As a result, an $\epsilon$-net argument shows that with probability $1-C(\sqrt{\frac{D}{D-d}})^{d(D-d)}\exp(-n_{\mathrm{out}}\alpha_{1}^{2}/8)$, (13) holds with $\beta_{2}$ in order of $\alpha_{0}\frac{\sigma_{\mathrm{in}}}{\sigma_{\mathrm{out}}}\sqrt{\frac{D}{(D-d)}}+1$, where the plus one term from the assumption $\beta_{2}\geq 1$.
Verification of (14): Similar to the proof of (13), WLOG we may assume that the inliers are sampled from $\mathcal{N}(0,\sigma_{\text{in}}\boldsymbol{P}_{L_{\star}}/d)$ and the outliers the outliers are sampled from $\mathcal{N}(0,\sigma_{\text{out}}\boldsymbol{I}/D)$.
Following a classic result in random matrix theory (Wainwright, 2019, Example 6.2), we have
$\Pr\left(\sigma_{1}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\boldsymbol{x}\boldsymbol{x}^{\top}\right)\leq\frac{\sigma_{\mathrm{out}}^{2}}{D}(\sqrt{n_{\mathrm{out}}}+\sqrt{D}+t)^{2}\right)\geq 1-\exp(-t^{2}/2).$
On the other hand,
$\min_{\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}\,:\,|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|}\sigma_{d}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\boldsymbol{x}\boldsymbol{x}^{\top}\right)=\min_{\|\boldsymbol{u}\|=1}\min_{\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}\,:\,|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|}|\boldsymbol{u}^{\top}\boldsymbol{x}|^{2}.$
Defining $\mathcal{X}_{\mathrm{in},\boldsymbol{u}}=\{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}:|\boldsymbol{x}^{\top}\boldsymbol{u}|\leq q_{\alpha_{0}}(\{|\boldsymbol{x}^{\top}\boldsymbol{u}|\}_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}}\}$,
$\sqrt{\min_{\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}\,:\,|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|}|\boldsymbol{u}^{\top}\boldsymbol{x}|^{2}}-\sqrt{\min_{\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}\,:\,|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|}|\boldsymbol{v}^{\top}\boldsymbol{x}|^{2}}=\sqrt{\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}},\boldsymbol{u}}|\boldsymbol{u}^{\top}\boldsymbol{x}|^{2}}-\sqrt{\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}},\boldsymbol{v}}|\boldsymbol{u}^{\top}\boldsymbol{x}|^{2}}$
$\leq$ $\sqrt{\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}},\boldsymbol{v}}|\boldsymbol{u}^{\top}\boldsymbol{x}|^{2}}-\sqrt{\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}},\boldsymbol{v}}|\boldsymbol{u}^{\top}\boldsymbol{x}|^{2}}\leq\|\boldsymbol{u}-\boldsymbol{v}\|\|\boldsymbol{X}_{\text{in},\boldsymbol{u}}\|\leq\|\boldsymbol{u}-\boldsymbol{v}\|\|\boldsymbol{X}_{\text{in}}\|,$
where $\boldsymbol{X}_{\text{in}}\in\mathbb{R}^{n_{\mathrm{in}}\times D}$ is the data matrix of inliers, with (Wainwright, 2019, Example 6.2)
$\Pr\left(\|\boldsymbol{X}_{\text{in}}\|\leq{\frac{\sigma_{\mathrm{in}}}{\sqrt{d}}}(\sqrt{n_{\mathrm{in}}}+\sqrt{d}+t)\right)\geq 1-\exp(-t^{2}/2)$
For any fixed $\boldsymbol{u}$, $\boldsymbol{x}\in\mathcal{X}_{\text{in}}$ is sampled from $N(0,1/d)$. As a result, Lemma 12(e) implies that for a fixed $\boldsymbol{u}$, $\min_{\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}\,:\,|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|}\frac{1}{n_{\mathrm{in}}}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in},\boldsymbol{u}}}|\boldsymbol{u}^{\top}\boldsymbol{x}|^{2}=O(\alpha_{0}^{3}\sigma_{\text{in}}^{2}/d)$ with probability at least $1-\exp(-n_{\mathrm{in}}\alpha_{0}^{2}/2)$. Applying a covering argument again over $\boldsymbol{u}\in S^{D-1}\cap L_{\star}$ with balls of radius $\epsilon=O(\alpha_{0}^{3}\sigma_{\text{in}}/d)$, we have that
$\Pr\left[\min_{\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}\,:\,|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|}\sigma_{d}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\boldsymbol{x}\boldsymbol{x}^{\top}\right)\geq c\frac{\alpha_{0}^{3}n_{\mathrm{in}}\sigma_{\mathrm{in}}^{2}}{d}\right]\geq 1-O\left(\left(\frac{d}{\alpha_{0}^{3}\sigma_{\text{in}}}\right)^{d}\right)\exp(-n_{\mathrm{in}}\alpha_{0}^{2}/2)$
Thus, (14) holds (and Assumption 3 holds) with probability at least $1-O\left(\left(\frac{d}{\alpha_{0}^{3}\sigma_{\text{in}}}\right)^{d}\right)\exp(-n_{\mathrm{in}}\alpha_{0}^{2}/2)-\exp(-n_{\mathrm{out}}/2)$ when
$\frac{n_{\mathrm{in}}}{n_{\mathrm{out}}}\gtrsim\beta\frac{d\sigma_{\mathrm{out}}^{2}}{D\alpha_{0}^{3}\sin(\theta_{0})\sigma_{\mathrm{in}}^{2}}.$
Summary: Recall that in our model we have $\theta_{0}=\pi/4$ and $\alpha_{0}=1/4$. Therefore, $\beta_{1}=O(1)$ and $\beta_{2}=O(\alpha_{0}\frac{\sigma_{\mathrm{in}}}{\sigma_{\mathrm{out}}}\sqrt{\frac{D}{(D-d)}}+1)$, and $\beta=10\beta_{1}\beta_{2}$, which implies that Assumption 3 holds with probability tending to $1$ as $n,D\rightarrow\infty$ when (24) holds.
Proof of Lemma 12.
(a) WLOG we may assume $\boldsymbol{\Sigma}=\operatorname{diag}(\sigma_{1},\cdots,\sigma_{D})$. Since $\chi^{2}$ is sub-exponential with parameter $(2,4)$ (Ghosh, 2021), by the property of sum of sub-exponential functions (Wainwright, 2019, Theorem 2.19 (a)), $\sum_{i=1}^{n}\|\boldsymbol{x}_{i}\|^{2}$ is also sub-exponential with parameter $(2n\mathrm{tr}(\boldsymbol{\Sigma}),4)$, and (Wainwright, 2019, Theorem 2.19 (b)) implies
$\Pr\left(|\sum_{i=1}^{n}\|X_{i}\|^{2}-\operatorname{\mathbb{E}}\sum_{i=1}^{n}\|X_{i}\|^{2}|\geq t\right)\leq e^{-\frac{tn}{8}},\,\,\text{for $t\geq\frac{\mathrm{tr}(\boldsymbol{\Sigma})}{2}$.}$
Recall $\operatorname{\mathbb{E}}\sum_{i=1}^{n}\|X_{i}\|^{2}=n\mathrm{tr}(\boldsymbol{\Sigma})$ and plug in $t=\mathrm{tr}(\boldsymbol{\Sigma})$, then
$\Pr\left((\sum_{i=1}^{n}\|X_{i}\|)^{2}\geq n\sqrt{2\mathrm{tr}(\boldsymbol{\Sigma})}\right)\leq\Pr\left(\sum_{i=1}^{n}\|X_{i}\|^{2}\geq 2n\mathrm{tr}(\boldsymbol{\Sigma})\right)\leq e^{-\frac{n\mathrm{tr}(\boldsymbol{\Sigma})}{8}}.$
(b) The first inequality follows from Hoeffding’s inequality applied to $I(x_{1}\leq q_{\alpha_{0}/2}(X)),\cdots,I(x_{n}\leq q_{\alpha_{0}/2}(X))$ (i.e., to the empirical cdf). The proof of the second inequality is similar.
(c) Applying (b), we only need to investigate the $\alpha_{0}/4$-th quantile of the population distribution of $\|\boldsymbol{x}_{i}\|$. We use the fact that for a positive random variable $X$, for any $t<0$, (the second inequality follows from the first inequality with a=1/ta=-1/t)
$\Pr(X\leq a)\leq e^{-ta}\operatorname{\mathbb{E}}[e^{tX}],\,\,\Pr(X\leq-1/t)\leq e\cdot\operatorname{\mathbb{E}}[e^{tX}].$ (70)
Note that the moment generating function of a $\chi^{2}_{1}$ distribution is $(1-2t)^{-1/2}$; so when $X=\|Z\|^{2}$ with $Z\sim N(0,\operatorname{diag}(\sigma_{1},\cdots,\sigma_{D}))$, is $\operatorname{\mathbb{E}}[e^{t\|Z\|^{2}}]=\prod_{i=1}^{D}(1-2\sigma_{i}t)^{-1/2}\leq(1-2\mathrm{tr}(\boldsymbol{\Sigma})t)^{-1/2}$ (the last inequality follows from $t\leq 0$). Let $t=\frac{1-16e^{2}/\alpha_{0}^{2}}{2\mathrm{tr}(\boldsymbol{\Sigma})}$, we have that the $\alpha_{0}/4$-th quantile of the population distribution of $\|\boldsymbol{x}_{i}\|$ is $\sqrt{-1/t}=\sqrt{\frac{2\mathrm{tr}(\boldsymbol{\Sigma})}{16e^{2}/\alpha_{0}^{2}-1}}\geq\frac{\alpha_{0}}{4e}\sqrt{2\mathrm{tr}(\boldsymbol{\Sigma})}.$
(d) Following part (b), it is sufficient to show that
$\Pr\left(\chi^{2}_{p}\geq(\sqrt{p}-2\log(\alpha_{0}/4))^{2})\right)\leq\alpha_{0}/4.$
This follows from the tail bound of that chi-squared distribution in Lemma 1 of Laurent and Massart (2000).
As for the second part, it is sufficient to show that
$\Pr\left(\chi^{2}_{p}\leq\left(1-\sqrt{\frac{4}{p}\log\frac{4}{\alpha_{0}}}\right)p\right)\leq\alpha_{0}/4.$
It follows from (Barvinok, 2005, Corollary 2.3) with setting $\epsilon=\sqrt{\frac{4}{p}\log\frac{4}{\alpha_{0}}}$, and note that $\epsilon\leq 1$ when $\alpha_{0}\leq 4\exp(-4/p)$.
(e) The first part follows from part (b), as well as the fact that the distribution of $N(0,1)$ has a density bounded above by $1/\sqrt{2\pi}$, so $\int_{-\alpha_{0}/2}^{\alpha_{0}/2}\varphi(x){\,\mathrm{d}}x\leq\alpha_{0}/\sqrt{2\pi}<\alpha_{0}/2$.
The second part follows from applying Hoeffding’s inequality to $X^{2}I(|X|\leq\alpha_{0}/2)$ where $X\sim N(0,1)$, which has valued bounded in $0,\alpha_{0}^{2}/4]$ and has expectation larger than $\frac{\alpha_{0}^{2}}{4}\cdot 0.35\alpha_{0}\geq 0.08\alpha_{0}^{3}$, where the factor $0.35$ follows from $\Pr(-\frac{\alpha_{0}}{2}\leq g_{i}\leq\frac{\alpha_{0}}{2})\geq\frac{\exp(-\alpha_{0}^{2}/8)}{\sqrt{2\pi}}\alpha_{0}\geq\frac{\exp(-1/8)}{\sqrt{2\pi}}\alpha_{0}\geq 0.35\alpha_{0}$.
6.4 Proof of Proposition 2
Proof.
We begin by showing that if condition (25) holds, then Assumption 2 is satisfied with high probability. Since outliers lie on the sphere, we have the trivial bound $S_{\mathrm{out}}\leq n_{\mathrm{out}}$. On the other hand, by Lemma 11, $S_{\mathrm{in}}\gtrsim Cn_{\mathrm{in}}\sigma_{\mathrm{in}}/\sqrt{d}$ with high probability. Combining these, we see that (25) is a sufficient condition for Assumption 2 to hold with high probability, for a fixed $\theta_{0}$.
For Assumption 3, we again assume that $\alpha_{0}=1/4$. Since the inliers follow a standard Gaussian distribution, by the verification of (12) step in the proof Proposition 1, (12) holds with high probability as in Proposition 1. As for (13), it is not needed as $\gamma>1-(1-\alpha_{0})\gamma_{\star}$ holds (see discussion after (13)). Thus it remains to verify (14). On the one hand, we again have the trivial bound as all outliers has magnitudes $1$:
$\sigma_{1}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\boldsymbol{x}\boldsymbol{x}^{\top}\right)\leq n_{\mathrm{out}}.$
On the other hand, for the inliers, following the proof of Proposition 1, we have the bound
$\min_{\mathcal{X}_{\mathrm{in}}^{\prime}\subseteq\mathcal{X}_{\mathrm{in}}\,:\,|\mathcal{X}_{\mathrm{in}}^{\prime}|\geq\alpha_{0}|\mathcal{X}_{\mathrm{in}}|}\sigma_{d}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\boldsymbol{x}\boldsymbol{x}^{\top}\right)\geq c\frac{\alpha_{0}^{3}n_{\mathrm{in}}\sigma_{\mathrm{in}}^{2}}{d}$
with high probability, and (14) then follows. ∎
6.5 Proof of Propositions 3 and 4
Proof of Proposition 3.
Let $a=\operatorname{dist}(0,[A])$, where $A]=[(L,\boldsymbol{m})]$, and denote the largest principal angle between $L$ and $L_{\star}$ by $\theta_{1}$. Then,
$\frac{1}{n}\sum_{\mathcal{X}_{\mathrm{in}}}\|\boldsymbol{Q}_{L}\boldsymbol{x}\|=\frac{1}{n}\sum_{\mathcal{X}_{\mathrm{in}}}\sin(\angle(\boldsymbol{x},L))\|\boldsymbol{x}\|\leq\frac{1}{n}\sum_{\mathcal{X}_{\mathrm{in}}}\theta_{1}\|\boldsymbol{x}\|\overset{n\to\infty}{\to}O\left(\frac{\alpha_{\mathrm{in}}\theta_{1}}{\sqrt{d}}\right).$
Thus,
$\frac{1}{n}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\operatorname{dist}(\boldsymbol{x},[A])$ $=\frac{1}{n}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\|\boldsymbol{Q}_{L}(\boldsymbol{x}-\boldsymbol{m})\|$ (71)
$\leq\frac{1}{n}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\|\boldsymbol{Q}_{L}\boldsymbol{x}\|+\frac{1}{n}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\|\boldsymbol{Q}_{L}\boldsymbol{m}\|$
$=O\left(\alpha_{\mathrm{in}}\left(\frac{\theta_{1}}{\sqrt{d}}+a\right)\right).$
On the other hand, for any $\boldsymbol{x}$, we have
$\operatorname{dist}(P_{A_{\star}}\boldsymbol{x},[A])\leq\operatorname{dist}(P_{A_{\star}}\boldsymbol{x},L)+a\leq O\left(\theta_{1}\|P_{A_{\star}}\boldsymbol{x}\|+a\right).$
As a result, the following inequality
$\frac{1}{\cos^{2}c_{0}}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\operatorname{dist}(\boldsymbol{x},A)\geq 4\cdot\frac{\pi}{2}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\operatorname{dist}(P_{A_{\star}}\boldsymbol{x},[A])$
holds when
$\frac{\alpha_{\mathrm{in}}}{\alpha_{\mathrm{out}}}\geq O\left(\max\left({\sqrt{d}\int_{\boldsymbol{x}\sim\mu_{out}}\|P_{A_{\star}}\boldsymbol{x}\|},\ 1\right)\right).$
Furthermore, applying (71) again, we also obtain the following inequality
$\frac{1}{\cos^{2}c_{0}}\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{in}}^{\prime}}\operatorname{dist}(\boldsymbol{x},A)\geq 4\cdot\theta_{1}(L,L_{\star})^{2}\left\|P_{L_{\star}^{\perp}}\left(\sum_{\boldsymbol{x}\in\mathcal{X}_{\mathrm{out}}}\boldsymbol{x}\boldsymbol{x}^{\top}\right)P_{L_{\star}^{\perp}}\right\|$
under the condition:
$\frac{\alpha_{\mathrm{in}}}{\alpha_{\mathrm{out}}}\geq\theta_{1}(L,L_{\star})^{2}\cos^{2}c_{0}\cdot\sqrt{d}\left\|\int_{\boldsymbol{x}\sim\mu_{out}}P_{L_{\star}^{\perp}}\boldsymbol{x}\boldsymbol{x}^{\top}P_{L_{\star}^{\perp}}\right\|,$
when $\theta_{1}(L,L_{\star})\leq c_{0}$.
Combining the estimations above, the proposition is proved.
Proof of Proposition 4.
Recall that the inliers are distributed as $\mathcal{N}(0,\boldsymbol{\Sigma}_{\mathrm{in}}/d)$, while the projected outliers $P_{A_{\star}}\boldsymbol{x}$ follow the distribution $\mathcal{N}(0,\boldsymbol{P}_{L_{\star}}\boldsymbol{\Sigma}_{\mathrm{out}}\boldsymbol{P}_{L_{\star}}/D)$.
Since the function $\operatorname{dist}(\boldsymbol{x},[A])$ is convex in $\boldsymbol{x}$, and using the assumption $\sigma_{d}(\boldsymbol{\Sigma}_{\mathrm{in}})\geq 1$, we obtain
$\int_{\boldsymbol{x}\sim\mu_{\mathrm{in}}}\operatorname{dist}(\boldsymbol{x},[A])\geq\int_{\boldsymbol{x}\sim\mu_{0}}\operatorname{dist}(\boldsymbol{x},[A]),$
where $\mu_{0}\sim\mathcal{N}(0,\boldsymbol{P}_{L_{\star}}/d)$.
Similarly, since $\sigma_{1}(\boldsymbol{\Sigma}_{\mathrm{out}})^{2}\leq D/d$, it follows that
$\boldsymbol{P}_{L_{\star}}\boldsymbol{\Sigma}_{\mathrm{out}}\boldsymbol{P}_{L_{\star}}/D\preceq\boldsymbol{P}_{L_{\star}}/d,$
and thus
$\int_{\boldsymbol{x}\sim\mu_{0}}\operatorname{dist}(\boldsymbol{x},[A])\geq\int_{\boldsymbol{x}\sim\mu_{\mathrm{out}}}\operatorname{dist}(P_{A_{\star}}\boldsymbol{x},[A]).$
Additionally, observe that
$\left\|\int_{\boldsymbol{x}\sim\mu_{\mathrm{out}}}P_{L_{\star}^{\perp}}\boldsymbol{x}\boldsymbol{x}^{\top}P_{L_{\star}^{\perp}}\,d\boldsymbol{x}\right\|\leq\frac{\|\boldsymbol{\Sigma}_{\mathrm{out}}\|^{2}}{D}\leq\frac{\sigma_{\mathrm{out}}^{2}}{D}.$
Combining this bound with the inlier estimation (71) and using the assumption $\theta_{1}\leq c_{0}$, the proposition follows. ∎
7 Conclusion
In this work we have revisited the use of IRLS algorithms for robust subspace recovery. In the case of linear subspaces, which is the FMS algorithm Lerman and Maunu (2018a), we have incorporated dynamic smoothing, which allows us to prove a global recovery result under a deterministic condition. We have extended the FMS algorithm to the affine setting, where we are able to show a local recovery result under a deterministic condition. These results represent significant theoretical contributions: this paper is the first to demonstrate global recovery for a nonconvex IRLS procedure and for IRLS over a Riemannian manifold, and it is the first to include a convergence analysis for an algorithm in affine robust subspace recovery. We include an example application of the (A)FMS algorithms to low-dimensional training of neural networks, where the found subspaces perform better than alternatives like PCA.
There are many open questions for future work. While our results are stable to small perturbations of the inliers, detailed noise analysis in robust subspace recovery is still an open question. In the case of affine subspace recovery, it would be interesting to show if a similar global convergence result holds as in the linear case or if there are some obstructions to such a result. Another avenue is to develop a deeper understanding of how robust subspaces have better generalization in the training of neural networks, and to leverage this in specialized subspace recovery algorithms for this setting.
8 Acknowledgements
G. Lerman was supported by the NSF under Award No. 2152766. T. Maunu was supported by the NSF under Award No. 2305315. T. Zhang was supported by the NSF under Award No. 2318926.